Adverbs in Focus

Regine Eckardt DYANA/Universität Tübingen

Abstract

My main aim is the study of predicate adverbs in German. It turned out, however, that a safe understanding of what is going on in sentences with predicate adverbs required to investigate a different thing first. For weak determiners, and indefinite NPs in special, it has been observed that in some cases they show what is called a partitive reading. Now the meaning of sentences with predicate adverbs as understood by speakers led me to suspect that not only a partitive interpretation of the NPs in question was involved, but that the readings could only be accounted for if referring to further accents in the sentence that have influence on the computation of restrictors and scopi of the quantifiers in question. This effect moreover arises not only with weak, but with any sort of determiners. The study of interaction between nominal quantifiers and focus is subject of the first part of the paper. In the second part, I come back to the original question of, how adverbs of manner behave in German, or rather, how NPs behave when moved around the adverbs. The preparatory work done in part 1 allows us to give clearer statements on what happens. In section 7, I want to show that interaction of focus with nominal quantifiers indeed is the cruical factor to safe adverbs in examples with negation from being unacceptable. This stands in contrast to the suspicion held up to now that ill-formedness or well-formedness of sentences with negation and predicate adverbs was a mere matter of checking the correct constellations at LF.

Contents

PART 1: Focus in interaction with nominal quantifiers 292

1 Introducing the data 292

2 Weak and strong determiners 294

3 Counting events: Washed cars and passing ships 299

4 A formal account 304

5 Comparison with Topic-Focus Theories 308

PART 2: An application in adverbial semantics 313

6 Adverbs of manner in German 313

7 Adverbs and Negation 322

8 Appendix 325

9 Bibliography 330

PART 1: Focus in interaction with nominal quantifiers

1 Introducing the data

The interaction of focus with adverbial quantifiers is a process widely discussed in the literature; it has been convincingly shown that focus indeed is not only a pragmatic process, but influences the truth value semantic of sentences. Less noted, and less understood, is the fact that also nominal quantifiers can interact with focus. My aim is to give a survey over the data, discuss various aspects of the phenomenon in more detail and give a formal account in the end. Also a comparison with a competing treatment will be given. Consider the following data

1.1 Max washed most/ cars with X-polish\. 1.2 Almost all/ tickets were sold at checker 4\. 1.3 Jan sliced no/ bagel carefully\.

where the ,,/“ is to be read as a raise accent, ,,\“ as a fall accent on the appropriate syllable. What should these sentences mean according to ordinary semantic, and what do we actaully understand them to say? 1.1 should be true in the world exactly if there is a set of (relevant) cars present most of which Max has washed - and moreover: washed with X-polish. For the rest of cars, we do not know. But actually if 1.1 is uttered in the intonation pattern indicated, we do well know about the rest of the cars: we understand that there is a set of cars washed by Max under discussion, and most of these washed cars have been washed with X-polish. It is easy to check that these two readings actually differ in truth value behaviour. The same observations can be made for sentence 1.2 and 1.3. The second example is understood not as saying that almost all tickets of some set of tickets were sold, but that allmost all of the actually sold tickets were sold at a special checker, namely checker 4. The third example does not get a different truth value behaviour under the indicated accenting, but different felicity conditions in terms of presuppositions: whereas nothing in a traditional representation of the meaning of the sentence would force us to do so, we nevertheless understand it as reporting that Jan has sliced bagels: and none of these bagel-slicings were done carefully.

When we compare the examples above with other cases of interaction of focus with quantification, we can easily observe that as always the cruical rule of the thumb governing the interpretive process can be given as: take the background of the sentence to give the restrictor argument of the quantifier, take the focussed material to go in the scope, and then: quantifiy along! This is, in a nutshell, both the essence of the treatment of ,,only“ as well as adverbial quantifiers with focus, as proposed in various papers by Rooth, Krifka, von Stechow, and others. We can now try and spell out this rule of the thumb in a first formalization of what is going on in 1.1 to 1.3. The following suggestion was made by Geilfuß[93]:

It is wellknown that nominal quantifiers always show a certain interaction with context. So, if I utter the sentence:

1.4 Every child had an ice cream.

I probably do not mean by this that all children in the entire world had an ice cream at the occasion, but that of a certain set of childern given in a context all had an ice cream. This context dependency has been given an explication by Dag Westerståhl (Westerståhl [85]), who proposed to use a parametrized version of nominal quantifiers Q_{C}(A, B) with Q_{C}(A, B) holds iff Q(A ∩ C, B) holds for the unparametrized version of the quantifier.

To go back to our example, this would give: [[ Every child had an ice cream ]] = ALL_{C}(λx.CHILD(x), λx.EAT −ICE(x)) where C was instantiated by the context. So, if f.i. C would be the set of all persons present at the party at Schroeder’s last Sunday, the sentence in that context would be represented as

ALL_{C}(λx.CHILD(x) ∩ λx.P ERSON − AT − SCHR(x), λx.EAT − ICE(x))

which actually says: All children at Schroeder’s had an ice cream. Now Geilfuss proposes to use exactly this parameter to compute the appropriate readings for examples like 2.1 - 3. Abstracting away for the moment from the problems that arise if we want to adopt the mechanism for arbitrary NPs to get their determiner interact with a focus in the sentence, it can be stated as follows: take the focus semantic value of the ,,rest of the sentence“ - without the NP - which will be a set of properties A, which all differ in the place indicated by the focus (= fall accent) and in that place exploit all relevant alternatives to the focussed item in the context under discussion. Now take ∪A which will yield a property, and instantiate the context parameter of the nominal quantifier that is to interact with the focus with this value ∪A. Thus the restrictor argument of the nominal quantifier will get cut down by the focus-background of the sentence and it turns out that in the above examples this will produce exactly the intended readings. Let us go through example 1.2 to see how this will work out. The quantifier to interact with focus is ALMOST − ALL. The ordinary interpretive process will give the meaning of the noun ,,tickets“ as the restrictor argument of this quantifier, and the VP ,,sold at checker 4“ as the scope. Now we are to compute the focus semantic value of the VP,

[[ sold at checker4_{f}]]_{f} =
{λx.∃y(SOLD(x, y) ∧ y = CHECKER1),
λx.∃y(SOLD(x, y) ∧ y = CHECKER2),
λx.∃y(SOLD(x, y) ∧ y = CHECKER3),
λx.∃y(SOLD(x, y) ∧ y = CHECKER4)}

where we assume that in the given context the relevant alternatives for tickets to be sold consist of the four checkers1 - 4. Now take the union of this set. This will be the property

λx.∃y(SOLD(x, y) ∧ (y = CHECKER1 ∨ y = CHECKER2
∨y = CHECKER3 ∨ y = CHECKER4))

As for the given context, tickets are taken to only be sold at one of these checkers, if at all, this is for the purpose equivalent to the property of being sold anywhere, i.e.: λx.SOLD(x). Now instantiate the context parameter of ALMOST − ALL with this property. This will give us as the meaning of 1.2:

ALMOST − ALL[λx.T ICKET (x) ∩ λx.SOLD(x),
λx.∃y(SOLD(x, y) ∧ y = CHECKER4)]

in perfect match with the paraphrase we gave in the beginning. So here we have got some impression of how a treatment of the examples could look like. The philosophy of the given formalization could be described as: focussing things means that other things are given by context, that could stand instead. Thus the parameter where nominal quantifiers show their sensitivity to context is well the place where a focal influence on the arguments of the quantifier can enter. Moreover the treatment fits the overall picture of focus in quantificational contexts insofar as it is just an adaption of the theory for focus and adverbs of quantification Rooth[92].

2 Weak and strong determiners

To start off, we shall try and vary the choice of determiners involved, and then see what happens. In special, it will be interesting to contrast the behaviour of weak and strong quantifiers under in these raise fall intonation patterns. (I will call these constructions FNQ constructions in the following; FNQ for focus with nominal quantifiers.). In order to be able to appreciate the following results, we will briefly recapitulate the basic story on weak and strong quantifiers as it can, for instance, be found in Partee[88]. The distinction between ,,strong“ and ,,weak“ quantifiers dates back to Milsark [74], who observed that in ,,there“-sentences in English not all quantifiers are equally acceptable. So whereas

3.1 There is a/ one mouse in the attic.

3.2 There were many/ few/ no/ some/ five mice in the attic.

all are good sentences of English, the following parallel constructions are bad:

3.3 *There are all/ most/ most of the/ the/ none of the/ half of the mice in the attic.

The determiners in 3.1 and 3.2 were called ,,weak“, and the ones in 3.3 ,,strong“ by Milsark. Apart from this empirical difference, it has also been observed early, that weak and strong determiners differ with respect to the nature of formal object their meaning most appropriately gets represented by. For strong determiners, the general format of object denoted seems to be a binary operator on sets, or generalized quantifier, Q(A,B), where the restrictor argument A (which corresponds roughly to the N coming with the determiner) is assumed to come along with a nonemptyness presupposition. So, a sentence like:

3.4 Every dog was excited.

is to be read as: ,,The set of dogs is non-empty, and is a subset of the set of excited beings“. The weak determiners in turn have a socalled ,,existential“ meaning. They do not presuppose, but report the non-emptiness of certain sets (and besides qualify the size of the sets, sometimes, like with ,,many“, ,,few“ and the cardinals ).

The next step in the study of weak and strong determiners was the observation that in certain constructions, weak determiners still seem to share the properties of the strong ones. This can be seen most clearly with the determiners ,,many“ and ,,few“, because here a real difference in truth conditional behavoiur takes place, whereas for the other weak determiners the shift from existential to quantificational reading only manifests in a change of presuppositional behaviour. So compare

3.5 Many children wanted to have ice cream after lunch.

3.6 Many cowboys entered the saloon.

Sentence 3.6 is most prominently understood to report that, from this point on, a (comparatively large) set of cowboys is there. One would most naturally represent this as

3.6′∃X(COW BOY S(X) ∧ ENT ER(X, SALOON) ∧ |X| > e)

where e is a contextually determined value for the standard size of groups of cowboys who enter saloons. In special, the cowboys we talk about are all cowboys we know of at this point. Sentence 3.5 in contrast means something else: Of the children present at lunch, some larger part also wanted to have ice cream afterwards.

3.5’ MANY (λx.CHILD(x), λx.W ANT S(x, ICE − CREAM)) with MANY (A, B) true iff |A ∩ B|/|A| > e where e is a contextually given average value (here: of children desiring ice-cream per children in total).

So in this example, the set of children talked about is not all children present, but just a subset of these children. The appropriate representation of the determiner is by a binary operator. the restrictor argument is presupposed to be non-empty. An extensive discussion of the data that support the suspect that ,,weak“ determiners actually are ambiguous between a ,,strong“ and a ,,weak“ reading is given by Partee (Partee[89]), such that we face the following situation:

strong  weak
A: Milsark  ^{∗}there…  ^{√}there… – sentences
every, most,  many, few, some, one …
half of, …
strong  weak
B: Partee quantificational: quantificational: existential
(summary)  Q(A, B)  Q(A, B)
in+outside VP in+outside VP in VP only
ILP+SLP  ILP+SLP  SLP only

At this point we can come back to our original question, namely whether any differences in behaviour can be observed between weak and strong quantifiers in FNQ constructions. First we can check that in a great many constructions, weak and strong determiners behave exactly parallel in that focus - in the manner indicated in section 1 - modifies the restrictor argument of the binary quantifier involved in the construction. Note that this means in special that all weak determiners come with their strong meaning (the binary operator reading), whenever they are to interact with focus. There have been several suggestions made as to how one could predict this, suggestions that either use some double focus constructions or are formulated in terms of topic and focus. We will examine those later, but for the moment just take it as a fact that interaction with focus forces a weak determiner to be interpreted as a generalized quantifier (this could be one effect of interpreting the raise accent). So, the following set of data present a coherent picture:

3.7 Alma has milked almost all/ cows on the meadow\.

3.8 Olga fed half/ of the cats with milk\. 3.9 Selma sliced no/ bagel carefully\ 3.10 Alma fed many/ cats with milk\. 3.11 Few/ students have read the paper carefully\. 3.12 One/ student had a coffee with milk\.

Nevertheless it seems that weak determiners can also enter in a different focus construction, with a focus very close to the determiner, which find no analogon with strong determiners:

3.13 Many/ scandinavians\ have won the nobel prize in literature 3.14 Few/ women\ were at the conference. 3.15 * Most/ Scandinavians\ have won the nobel prize in literature. 3.16 * Every/ semanticist\ was at the conference.

How do we understand these FNQ constructions? Sentence 3.13 is to be paraphrased as: ,,Many of the winner of the nobel prize in literature came from: Scandinavia“ and in this paraphrase has exactly the truth conditions predicted by the ,,strong“ interpretation of ,,many“ with the unfocussed part as restrictor and the focussed as scope: With A = Winners of the nobel prize in literature and B = Scandinavians, the paraphrase says: ,, |A ∩ B|/|A| has a value greater than expected.“ This is something different from both, - saying that many people in Scandinavia are winners of a nobel prize (MANY (A, B) with A = Scandinavians, B = Nobel prize winners ) and

- saying that many of the scandinavian winners of a nobel prize actually did in in literature (MANY (A, B) with A = scandinavian winners of nobel prizes, and B = scandinavian winners of the nobel prize in literature ). These are the two most sensible proportional readings of the sentence to be produced, where the noun ,,scandinavians“ has import in the restrictor argument of the operator. Both paraphrases are not what we understand the sentence to mean. Moreover it actually is the proportional reding that is in play, because again, these ,,many“ nobel prize winners from scandinavia are not all nobel prize winners under discussion (like in the cowboy example), but are clearly compared in size to the overall set of winners. As it is always difficult to disambiguate readings with items that are as vague in meaning as ,,many“ and,,few“ are, I want to give a further example to exclude that we have to deal with some existential claim, together with certain focus effects.

3.17.a Last saturday I was at a party of anthroposophes.

b. Many/ women\ were among the drunk guests. b’. Few/ [drunk women\]_{f} were among the guests.

It seems to me that in the same situation both b. and b’. can be a correct description of the situation at the party (although not in sequence, because discourse topic and accents in focus constructions interact). Imagine that Anthroposophes are sober people who do not tend to excessively consume alcool even at festivous occasions. So the percentage of drunk people, and also then drunk women, compared to the overall number of guests was low. This is what sentence 3.17.b’ would report. Nevertheless the women Anthroposophes still lead a more frustrating life than their male companions, such that the percentage of drunk women among the drunk guests was higher than to expect - which is reported by 3.17.b . (The overall generalization would then be: Anthroposophes seldom drink, but if they do, they are often women. I want to emphasize that I never actaully have found such a pattern.) What would the existential reading give us for these sentences? In both cases it comes down to saying something about the size of the set of drunk female guests at the party, and it still would have to be explained how in one and the same situation a mere change of accent pattern could enforce the same set once to be smaller than predicted, and once greater than predicted. In comparison to the weak determiner examples, the parallel sentences with strong ones I gave have an asterisk (3.15 and 3.16). This does not mean that they are unacceptable, but just shall indicate that the accents given are not interpreted as to interact with the determiners, but rather interpreted as contrastive, or slip of the tongue, or whatever. So for instance, 3.16 does not mean that everybody at the conference was a semanticist, but, if something at all, that not everyone from a given goup of phonologists, or syntacticians, no, but from a given group of semanticists was at the conference (whoever else). So, to summarize what we have found up to now, indeed there are differences between weak and strong determiners with respect to FNQ constructions. For strong determiners, a focus on the noun coming in the same NP as the determiner does not interact with the quantification, whereas for weak determiners this pattern in principle is allowed. This has the interesting consequence that for weak determiners (in their ,,strong“, i.e. quantificational reading) the computation of the arguments of the operater is focusdriven, and not syntaxdriven as normally assumed for any sort of nominal quantifier: we saw that focus can even force the noun coming with the determiner to have semantic import in the scope, not in the restrictor of the quantifier. This in itself is interesting. But the full picture we find is even more complex than that, but in the end we will find that matters do not behave completely arbitrarily. Herburger has made the observation (Herburger[92]) that the possibility to interpret narrow raise fall pattern like in 3.13, 3.14 even for weak determiners covaries with two other tests that originally were thought to distinguish between strong and weak quantifiers: first, the position of the determiner in or outside VP (especially in German), and second the combination with stage level or individual level predicates. The cruical data are the following: It is known (f.i. from the work of Diesing, Diesing[92]) that existential readings for weak determiners only are good, or preferred, in positions in the sentence within VP (this is a part of Diesings mapping hypothesis). In German it is possible to localize the VP boundary by certain particles, like ,,ja doch,,, apart from various adverbs. Strong quantifiers (both as interpretations of weak or strong determiners) in turn are good both stemming from within or outside VP (at surface structure) which is thought to be due to the fact that at LF, they get rosen anyway. Now, the narrow FNQ pattern for weak determiners are good within, but not good outside VP:

3.18 *...weil viele/ Schwalben\ ja doch in den Sueden fliegen.

(because) many swallows  PRT to the south  fly

3.19 ... weil ja doch viele/ Schwalben\ in den Sueden fliegen.

(because) PRT  many swallows  to the south  fly

Whereas 3.18 evokes the same contrastive interpretation of the focus accent as the strong quantifier examples, 3.19 has the predicted reading where focus interacts with quantification: many of those who fly south are swallows. Second, Herburger diagnoses a contrast in acceptability of narrow hat focus on weak determiners, according to whether they are combined with individual level predicates or stage level predicates:

3.20 * Few/ boys\ are intelligent.

3.21 Few/ boys\ are in the soccer team.

Again, it has been held that individual level predicates (“be intelligent,,) are not acceptable with weak determiners in their existential readings, whereas stage level predicates (“be available,,, “be in the soccer team,,) are good. Strong quantifiers were observed to be good in both constructions. Now we are in a puzzling situation: Not only are the “weak determiners,, ambiguous between a weak (existential) and a strong (quantificational) reading, but in turn in the quantificational case there is a further subdivision between acceptable narrow raise fall pattern and uninterpretable narrow narrow raise fall pattern, which covaries with the previous weak/strong distinction.

strong  weak
A: Milsark  ^{∗}there…  ^{√}there… – sentences
every, most,  many, few, some, one …
half of, …
strong  weak
B: Partee quantificational:  quantificational:  existential
(summary)  Q(A, B)  Q(A, B)
in+outside VP  in+outside VP  in VP only
ILP+SLP  ILP+SLP  SLP only
C: with FNQ  quant  quant  quant  exist
in+out VP  in+out VP in VP  in VP
ILP  ILP  SLP  SLP
no narrow  no narrow ^{√}narrow  –?–
raise-fall  raise-fall raise-fall
strong^{∗}  weak^{∗}

It could be suggested here that a new classification of determiner readings in weak* and strong* is in place, where strong* consists of the two left rows of the table, and weak* of the two right ones. Is there any semantic criterion to support this distinction? In terms of quantification and presupposition the picture has become unclear here: Neither is it in the strong* case that always and automatically the content of the noun coming with the determiner is identical with the restrictor of the quantifier and this set in turn is identical to the set presupposed, - nor is it in the weak* case that generally nothing is presupposed, because we also find quantifiers among the weak* readings of determiners - but there, in turn, it is not allowed at all to take the content of the noun to go into the restrictor as usual, but funnily it is pressed into the scope of the cruical quantifier. I want to suggest that the weak*/strong* distinction of determiners is one to be expressed in terms of presupposing the content of the noun in the NP:

(A) A weak* determiner may never presuppose the content of N in its NP [ Det N ]_{NP} is only good if [[N]] still can be ∅ in the context. A weak* determiner stays in VP at LF. (B) A strong* determiner must presuppose the content of N in its NP: [ Det N ]_{NP} is only good if [[N]] is known to be nonempty. A strong* quantifier gets rosen at LF.

In case (A), this results in either reading Det as existential, where no presuppositions arise anyway, or else interpret Det well as binary operator which presupposes the content of its restrictor argument - but to force [[N]] to go into the scope of this operator and thus not be presupposed. In case (B), Det is in any case interpreted as quantificational. Either no focus is in play, then [[N]] is the restrictor argument, or else, Det may not c-command a focus. This namely, results in the restrictor argument A - even if exclusively computed by considering focus structure! - to be a subset of [[N]], such that, if A is presupposed to be nonempty, automatically [[N]] also must be nonempty, and the presupposition requirement placed by syntactic structure is met. Thus we can restate the influence syntax has on the interpretation of determiners and their arguments in terms of presupposition requirements. Moreover we can postulate a uniform focus driven interpretation mechanism for FNQ constructions instead of being forced to switch theories for weak and strong determiners.

At this point we have both learned something about determiners, and focus interacting with them. We have learned that focus suggests a new partition of determiner readings in weak* and strong*, and we have learned that in this new classification interaction with focus can be viewed as a mechanism as independent of ordinary syntax driven interpretations for NPs as focus with adverbial quantifiers is.

3 Counting events: Washed cars and passing ships

Let us go back to Ludwig and his car washings.

3.1 Ludwig washed most/ cars with X-polish\.

To get a more systematic picture of what it is we’re counting, let me scetch what the available options are. First: The re-quantificational reading. We have a set of cars Ludwig has washed, maybe several times. The re- quantificational reading would say something like: “Most of the cars Ludwig has washed at some time, he has also at some (possibly other) time washed by the use of Xpolish.,, It would be true in a situation where Ludwig has five cars at hand, C1 - C5, and has washed:

C1 - w1: with X  C2 - w1: with X  C3 - w1: without C4 - w1: without C5 - w1: without
- w2: without X  - w2: without  - w2: without  - w2: with X  - w2: without

- w3: without X - w3: with X

So, there were 8 washings without X-polish, 4 washings with X-polish, and the sentence would count as true. This is the reading that is obtained by the formalism given in section 1. Second: The event-counting reading. What the quantifier really quantifies over are the events under consideration - which formally is the same as counting pairs of events and objects, as long as no delicate plural ontology is in play. The sentence is then paraphrased as: “Most of the washings of cars by Ludwig were done with X-polish.,, In the scene given above, this reading would then be false. On the other hand, it would be true if the world looked the following:

C1 - w1: with X C2 - w1: without C3 - w1: without C4 - w1: without C5 - w1 : without - w2: with X - w2: with X

- w3: with X - w3: with X

So, there are 5 washings done with X-polish, and 4 done without. On the other hand, only two of the five cars were ever washed with X-polish, such that the first reading would be false in this scenario. This is a reading that we would formally produce if we decided to quantify unselectively over pairs of individuals and events, i.e. washings and cars in the example. Third: The one-one-reading. The whole problem is circumvented by claiming that a sentence like 5.1 is only uttered in situations where it is clear that Ludwig has washed each of the cars under consideration exactly once. Then the two readings above coincide and are moreover equivalent to quantification over cars only:

C1 - w1: with X C2 - w2: without C3 - w3: with X C4 - w4: without C5 - w5: with X

This reading would have to be paraphrased by: “There are washings of cars by Ludwig under discussion, and no matter how often the cars were washed in total, we are only interested in one washing per car and quantify over these. This accounts for the intuition that we talk about the same washings in the restrictor and scope (which is given up in the re-quantificational formalization in 1) and nevertheless doesn’t count events (like in 2). Note however that this ,,reading“ is logically indistinguishable from reading 1, because once we have established that reading 1 is true in a model, we can always claim that it were exactly the events that were chosen in the scope (,,there also is a washing that was done with X-polish“) that were under discussion and will make the one-one-reading true. So this third paraphrase is not really a reading, but unfortunately nevertheless is the most appropriate paraphrase for the examples.

The distinction between readings where events are counted and readings where objects are counted can also be found in other examples and has been investigated by M. Krifka in the paper ,,4000 ships passed through the lock“. He makes a suggestion on how to treat these sentences, and as a by-product of his theory, also an account for FNQ- cases was given. So, how does his formalism work? The original question was how to model the readings of the sentence

3.2 4000 ships passed the lock.

3.2 can be read as telling something about 4000 actual ships, namely that each of them passed the lock, but it can also be read as counting the passings through the lock, no matter how many actual ships exist - there might be much less, and some of them passed several times. So, Krifka develops a way to glue together passing events and count their size ,,in ships“, such that the second (event counting) reading would be given as: ,,there was an event of passing the lock, which was 4000 ships big“ where the measure 4000 ships does not imply that 4000 actual ships exist, but only that there exist smaller events of actual ships passing that are subevents of the big one, and that if you add up all ,,ship“ that occurs in such smaller events, you come up to a total of 4000. The first, individual counting reading is given in the traditional way. This makes sense especially in examples like

3.3 Most ships passed the lock

where not all ships were involved in a passing at all, and thus a restriction to counting events alone would miss the point. What happens if focus enters the picture?

3.4 Most/ ships passed the lock at night\.

Here again, we can diagnose that the sentence can be read in two ways: either as telling that most of a set of actual ships that passed the lock - passed the lock at night; or else as reporting that most passings happened at night, no matter if any ship passed twice, or three times, or... . So it seems that FNQ and event counting readings are two phenomena that stand orthogonal to each other. Nevertheless some dependency should be noted: whereas in the cases without interacting focus, quantification over individuals was necessary as an independend device, in examples like 3.4 all ships in question are ships which have taken part in some event, namely a passing, such that even when individual ships are to be counted, we can instead quantify over events and say:

,,There was a global NONITERATIVE event of passings of shipsize n, with a subevent of passings of shipsize m, such that m/n > 1/2.“ The predicate NONITERATIVE tells us that no two of the atomic subevents of passing were done by the same ship, such that indeed there were as many ships as passings, such that what we count are indeed individuals. And maybe even more. The paraphrase given above will be true if the world looks like this:

Applied to the washed cars example (3.1), this corresponds to what was called there the requantificational reading. But by choosing the noniterative event once and for all and then saying that of these once chosen events, most were done in a special manner, we are more in the spirit of the one-one paraphrase than with the representation where we quantify over individuals and only when the focussed predicate is interpreted chose the appropriate events per individual (= washings, passings) for which the focussed predicate holds also (= being a washing with X-polish, being a passing at night). So at this point, matters seem to suggest that in FNQ constructions quantification over events in all cases is not only a possible but even the more appropriate option: when focus interacts with nominal quantifiers, let the nominal quantifier bind not the individual, but the event parameter of the sentence, and then reconstruct its restrictor and scope according to the focus structure of the sentence (as suggested by the previous section). But before commiting to this strategy, let us check somee more examples.

When events and individuals fall apart. Up to now the only mismatch between individuals and events arose because one individual took part in several events of the type described. But there are also cases where we have less events than idividuals, namely when collective actions take place: liftings of the concert grand, or hirings of a secretary, etc. So, take a situation where five people lift a piano and we are told:

3.5 Most/ helpers lifted the piano [with a rope\]_{f}.

Clearly what we count in this example are the persons taking part in the lift, and not the event in question, which, as there only being one, cannot be sensibly quantified over with ,,most“. So either we take this example as evidence that the proposal to quantify over events instead of objects was too rash, or we have to explain what else could be going on here. We might claim that, as soon as the focus is to be integrated, we have to reinterpret the verb ,,lift“ so as to say ,,take part in the lifting“ rather than ,,lift the piano“. For these partial liftings we then get several events of ,,taking part in lifting“, one per individual, and now can go and quantify. Anyway, if this reinterpretation of the collective verb was due to the fact that FNQ takes place then it would at least be questionnable why we, although insisting in quantification over events, are nevertheless at least counting individuals insofar as a condition ,,make sure that individuals don’t outnumber events“ must be met. Still, we could reinterpret the verb for mere logical reasons, so because it doesn’t make sense to say that one event was ,,most“. Let us complicate the example one step further so as to exclude this.

Imagine that in a TV show, there are several players, A, B, C, D, E, F, and that they form varying teams that all have the task to lift a piano. Imagine further (to support an event counting reading) that we take the perspective of the woman who has to stop how long each team can lift the piano: so she is sitting in a separate room where the piano stands, and the teams come in, lift the piano, and go out again. The following lifts have taken place:

Now what can this sentence report?

3.7 Most/ players have lifted the piano [with a rope\]_{f}.

I think, it can either report that most of the players A - F in general had the strategy to lift the piano with a rope whenever they had to, which would be quasi a universal quantification over events of lifting for each player. Or, it can say that (from the perspective of the woman with the stop watch) it was so that of the persons entering the room and lifting pianos, most did the lift with the help of a rope - which means that we count pairs of lifts and lifters, or maybe participations in liftings.

But it can by no means say that most of the overall lifts of the piano were done ,,with a rope“, neither this being true if at least one player used a rope in the lifting, nor if all lifters used a rope, so that we would quantify universally over lifters - which would be the dual reading to the first one given, where we universally quantify over lifts for each player.

I take that these data strongly suggest that we should quantify over individuals in FNQ sentences to get the individual counting readings, and over pairs of individuals and events to get the event related readings. I do not want to go into the question of how appropriate it is in general to quantify over events to get the socalled event related readings. Note however that the main competing suggestion to treat these examples, namely quantification over stages of individuals, would also bring the correct results in our cases. (This can easily be seen by noting that there is a one-to-one correspondence between pairs of individuals and events, and stages of individuals during the time the events take place. As in our cases, only one type of event is in debate, we will not get trouble for the reason that various events might take place at the same time.)

What about the one-one reading then, which was originally the reason to take up the whole discussion? Remember that we wanted to account for the paraphrase of 3.1 which went like: ,,For most of the cars Ludwig washed, THE washing was done with X-polish“. We noted that the definiteness of the washings of cars by Ludwig seemed to be due to the fact that only certain washings of cars by Ludwig were under debate at all (namely one per car). I would like to suggest that the notion of ,,car washings by Ludwig that are at debate“ in itself is rather something that relates to the discourse context, or topic of the utterance. In any case it seems an effect that should be accounted for by different factors, and studied in independence. These examples lead one to favour a treatment of E-type pronouns (and related cases as these here) more along the lines of Heim [90] than by brute unselective binding. I will not pursue this issue any further here.

Up to now, we have not lokked at any examples where more than one quantificational NP was involved. Part of the investigations in the second half of the paper is dedicated to the question of what happens (in German) if two indefinite NPs are in the sentence. The perspective taken there is somewhat different , however, so that I want to give at least a survey of what might happen at this point.

Unproblematic are the cases where further NPs (indefinite or quantificational) are in focus. So we can say:

3.8 Half/ of the girls gave [ most of their frogs \]_{f} to Santa Claus.

3.9 Most/ of the boys gave [five \]_{f} of the aunts a kiss.

where we contrast “ most of their frogs“ with “7 toy cars“, “a chocolate bar“, etc. , or “five of the aunts“ with “three of the aunts“ or “seven of the aunts“. Of course, when focus on or within the second quantificational or indefinite NP is to interact with the nominal quantifier in the first quantificational NP, this can only be done if the first NP has scope over the second. Or rather: technically, the scope relation which is the other way round will amount to having a trivial focus semantic value [[ X ]]_{f} = {[[ X ]]_{o}} for the quantifier to interact with. In almost all cases this would result in a necessarily false formula of the form Q([[ X ]]_{o}, [[ X ]]_{o}). Empirically we do not get a reading of (any of) these examples which is necessarily false. So it does not make sense to assume that this way to compute the sentence is a life option at all. The scope relation “First (= raise accented) NP has scope over second NP (in focus) matches the linear order, too. So these constellations exhibit a coherent and clearly tractable behaviour.

Intuitions about acceptability and interpretation get shaken heavily, however, when it comes to sentences with NPs in the range of an FNQ construction but in focus-background. For German we will see later that the entire complex of word order phenomena, intonation regularities, genericity etc. is in play. Let us look at some English examples:

3.10 Half/ of the Gauls have beaten most Romans ardently \ 3.11 Half/ of the Gauls have ardently\ beaten most Romans 3.12 Few/ farmers have milked most cows carefully\ 3.13 Few farmers have carefully\ milked most cows

Now according to educated prejudice we expect these examples to work along the following lines:

3.10a Half of those Gauls who did < focus-background >, did it < focus> 3.12a Few of those farmers who did < focus-background >, did it <

focus >.

(And the same for 3.11 and 3.13). So this would be spelled out as:

3.10b Half of those Gauls who did < beat most Romans in e_{1} > did

it(=e_{1}) ardently. 3.12b Few of those farmers who did < milk most cows in e_{2} > did it

(=e_{2}) carefully.

Two things are to be observed. First, we have a tendency in 3.10 and 3.12 to understand the second quantificational NP as raise-accented, too. With no exact idea about how to treat multiple raise accents (but maybe see the suggestions made in section 5) it nevertheless seems to me that this intonation pattern leads one to understand:

3.10c “ Gauls beat Romans, and: half of the Gauls most of their roman

enemies: ardently“ 3.12c “Farmers milk cows, and: few of the farmers most of their cows:

carefully.“

Although these readings are interesting in themselves, they nevertheless are not what we are out for. So we must take care not to read along these lines (this is excluded in 3.11 and 3.13 anyway).

Second, if we thus exclude this mis-read, we get an idea of why maybe the sentences under discussion leave us with a somewhat vague idea of what they report. Literally they should say:

“There are Gauls for each of which there is an event e of BEATING-MOST-ROMANS. For half of these Gauls, the event is done AR-DENTLY.“

“There are farmers for each of which there is an event e of MELKING-MOST-COWS. For few of them, the event is done CAREFULLY.

Now, I find it hard to imagine that such complex events are referred to. Other informants seemed to have no doubts, however, that a sentence like

3.14 Every Gaul beats most Romans.

not only can report the existence of one global event e where

∀x(GAUL(x); MOST y(ROMAN(y); BEAT (x, y, e)))

holds, or the existence of a lot of atomic subevents e′ with

GAUL(x) ∧ ROMAN(y) ∧ BEAT (x, y, e‘)

but that we can also subdivide events along the participants’ axis:

the event e_{1} of Gaul 1 beating every Roman he beats the event e_{2} of Gaul 2 beating every Roman he beats the event e_{3} of Gaul 3 ...

For certain examples it is quite easy to have a concept of such events, especially with existential NPs:

Most/ guests shyly\ knocked three times at the door

For others I admitt I have difficulties to understand the in this way, like for instance

Most/ guests boredly\ watched most holiday slides.

It is an open question in event semantics what happens between global events and their atomic subparts. Our examples again provide new data and in turn will gain from every insight on this topic.

4 A formal account

I now want to give a formal account for interaction of nominal quantifiers with focus which takes care for the observations made above. The idea is, to take over a treatment for adverbial quantification with focus and adopt it for the nominal case. However, as I have argued that no unselective binding is in play, I will refrain from using a dynamic semantic formalism. As to account for the foci, I want to use alternative semantics for focus, as given in Rooth[85]. So the fall accents in the examples will be treated as focus features. In contrast to the final version of Rooth[85], however, we want to use focus semantic values at any point in the evaluation of (parts of) the sentence. Whereas the adverbial focus sensitive operators are all fine with focus semantic values on a propositional level, the nominal quantifiers, naturally, not only need the focus semantic values for properties to be fed with, but also operate on parts of the sentence semantics where other NP arguments still are missing, f.i. an object determiner is to interact with focus and has to have scope under the subject NP. Here is maybe the appropriate place to say a word on relative scopes. One thing I am decidedly sure about is, that if a nominal quantifier is to interact with focus in the way investigated here, it must have the focussed element in its scope. This statement may seem trivial, but isn’t in the light of further examples with raise- and fall-accents, to which I want to give a short introduction in the next section.

A question much more difficult to answer is the following: may we assume that the quantified NPs may undergo quantifier raising at LF in full freeness, as long as the relative position of the raise accented NP and the focussed element is not disturbed? For the german parallel examples, I would dare to claim that according to my intuitions about the data, relative quantifier scopes always follow the ouvertly visible order of the noun phrases. However, in german it is possible anyway to shift the NPs such as to match the ouvert word order with almost any order of scopi. So instead of assuming that fixing scopi is an LF matter, it might be that in German, it is generally better to chose the appropriate ouvert order of NPs (but see Frey[92] on this question). In English I am uncertain about how good it would be to read:

4.1 Almost all/ farmers have tried most chinese dishes reluctantly\.

as saying:

4.2 For most chinese dishes: almost all of the farmers who have tried them, they did it reluctantly.

Or, to give an example which works the other way round, can one understand:

4.3 Almost all farmers have tried most/ chinese dishes reluctantly\.

as telling:

4.4 For most of the chinese dishes such that almost all farmers have tried them, holds, that they did it reluctantly.

What makes matters worse in these cases is, that we are in the range of the twoquantifier-cases of section 3, where we have already seen that it is difficult to make sense of examples where a second quantifier is in the background part of the sentence. So, to interpret 4.3 as to mean something like 4.4 will come down to interpret a comparatively clear original sentence in a way that is already puzzling if in match with ouvert word order. So, for reason still not clear, it is not probable that the shift from 4.3 to 4.4 occurs in practice. For 4.1 and 4.2 it would be the other way round: by changing the scope order we come to a structure that is easier to understand than the original one. So are we then tempted to read 4.1 as 4.2, at least in some cases? I admitt that I do not know at the moment. It would be interesting to find clear examples, or speakers with decided intuitions about the examples in question. For the time being, I will, at least for English, take the view that movement of quantifiers in LF is in principle possible in our cases; but maybe the readings just never are convincing.

A semantic formalism: Semantic values for sentences are computed in an extensional manner with respect to models that consist of D_{e} = set of individuals D_{E} = set of events (which could also be introduced as a sort of individuals; we will get too much functional objects here which we will never need to use, but anyway) D_{t} = 0, 1 D_{<}a, b >= D_{b}^{D_{a}} the set of functions from D_{a} to D_{b} The interpretation of lexical items is given by an interpretation function [[ . ]] according to the following rules 1. For nouns N, [[ N ]]_{o} is an object N’ in D_{<e,t>}

2. For intransitive verbs V, [[ V ]]_{o} is an object V’ in D_{<e,<E,t>>}

3. For transitive verbs V, [[ V ]]_{o} is an object V’ in D_{<e,<e,<E,t>>>}

4. For adverbs A, [[ A ]]_{o} is an object A’ in D_{<e,<E,t>>}

5. For traces [[ t_{i}]]_{o} = x_{i} , a variable of type e with index i

6. For determiners, we have the respective generalized quantifiers as their values, so [[a]]_{o} = the function SOME in D_{<<et>,<et>,t>>} that maps two sets, A and B onto the value 1 in D_{t} iff A ∩ B ≠ ∅ [[every]]_{o} = the function EV ERY in D_{<<et>,<et>,t>>} that maps two sets, A and B onto the value 1 in D_{t} iff A is a subset of B; [[most]]_{o} = the function MOST in D_{<<et>,<et>,t>>} that maps two sets, A and B onto the value 1 in D_{t} iff |A ∩ B| > |A\B| [[many_{1}]]_{o} = the contextually dependent function MANY in D_{<<et>,<et>,t>>} that maps two sets, A and B onto the value 1 in D_{t} iff |A ∩ B|/|A| > e, where e is the value to be expected for this ratio; [[two]]_{o} = the function T W O in D_{<<et>,<et>,t>>} that maps two sets, A and B onto the value 1 in D_{t} iff |A ∩ B| > 2; etc...

7. For the event quantifier O_{E}, [[OE]]_{o} = the one place existential quantifier in D_{<E,t>,t>}

Interpretation of sentences at LF: I take the LF structure of a sentence to come up in the following shape: The quantificational NPs are raised to S by leaving a trace behind they are coindexed with. Possibly one will restrict the raising so as to match the ouvert order of scopi. The meaning of the verb can thus be combined with semantic objects that all are of type e (proper names, pronouns and variables) and stand in the correct position to fill the right argument place. In a similar way, I want to take account for the event parameter of the verb. The idea is that we have an event operator OE which functions as an existential quantifier for the event place of the verb. The base position of this operator probably is somewhere at the VP boundary, and I assume, to be precise, that it is over the position where predicate adverbs are situated. (There is no point in that. Either decision one takes influences the spellout of the interpretation rule for predicate adverbs; but both are technically possible.) I want to assume that it is also raised at LF to an appropriate position thus adopting the way to get the nominal arguments of the verb filled for the event place. The combination of semantic values for daughter nodes to give the value for the mother node are as follows:

1. [[[NP_{i} V ]_{v}1]]_{o} = [[ V ]]_{o} ( [[ NP_{i}]]_{o} )

2. [[[NP_{i} X ]_{v}i]]_{o} = [[ X ]]_{o} ( [[ NP_{i}]]_{o} ) i.e within VP, any NP meaning is combined with its neighbour predicate by functional application of the latter to the former.

3. [[ [ PAdv X ]_{v}i]]_{o} = λx_{e}λz.( [[ PAdv ]]_{o}(z)(x_{e})∧ [[ X ]]_{o}(x_{e})(z))

4. [[ [ DET N ]_{NP_{i}}]]_{o} = [[ DET ]]_{o} ( [[ N ]]_{o} )

5. [[[NP_{i} S ]_{s}]]_{o} = [[ NP_{i}]]_{o}(λx_{i} [[ S ]]_{o} )

6. [[[OE S ]_{s}]]_{o} = [[ OE]]_{o}(λx_{e} [[ S ]]_{o})

Focus semantic values:

1. For any part of a sentence A with semantic value [[ A ]]_{o}, if A itself carries the focus feature, then [[ A_{f}]]_{f} = { x — x is a contextually given alternative to A of the same semantic type as [[ A ]]_{o} }

2. [[ A ]]_{f} = { [[ A ]]_{o} } else.

3. The combination of focus semantic values is defined as: [[ A B ]]_{f} = { a — there is x in [[ A ]]_{f} and y in [[ B ]]_{f} and a is the semantic combination of x and y with the operation one would use to compute [[ [ A B] ]]_{o} }

Special rules for FNQ computing: Here I want to give rules to handle the raise accent on the nominal quantifier. As we have seen, there is reason to claim that this nominal quantifier interacts with focus like an adverbial quantifier. I want to assume that the nominal quantifier leaves the NP it comes with. The NP without the quantifier that rules the combination of the noun with the rest of the sentence is interpreted as if there was a relative pronoun instead, to intersect the property given by the noun with the property given by the rest. I assume that a determiner that carries the feature DET_{r} as a reflex of the raise accent firstly: is adjoined to the left of its NP and secondly: is interpreted differently from an ordinary determiner. So we get:

1. [ [ DET_{r} N ]_{N}P X ]_{s} ⟶ [ DET_{r} [ [ - N ]_{N}P X ]_{s}]_{s} as an adjunction rule. and the two further interpretation rules:

2. [[ [ - N ]_{N}P i X ] ]]_{o} = [[ N ]]_{o} ∩ λx_{i}[[X]]_{o}

3. [[ [ DET_{r} X ]_{s}]]_{o} = [[ DET ]]_{o}(∪ [[ X ]]_{f} , [[ X ]]_{o} )

To see what happens, let us go through an example.

,,Max washed most/ cars care\fully“

will come with an LF that looks roughly like:

I have given the nodes in the tree numbers, for easier reference. We start interpreting at V_{1} which will give: for [[ washed ]]_{o} = λxλyλz.wash′(x, y, z) in D_{<e,<E,<e,t>>>} The object NP comes as [[ t_{i}]]_{o} = x_{i} and is combined with the verb to give [[VP_{3}]]_{o} = λyλz.wash′(x_{i}, y, z) The adverb is interpreted as λuλw.careful′(u, w) in D_{<e,<E,t>>} and combined to give

[[VP_{5}]]_{o} = λx_{e}λz.(λuλw.careful′(u, w)(z)(x_{e}) ∧ λyλz.wash′(x_{i}, y, z)(x_{e})(z))
= λx_{e}λz.(careful′(z, x_{e}) ∧ wash′(x_{i}, x_{e}, z))

Now the event parameter and the subject are filled in to give For [[S_{7}]]_{o} = careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX) Next, the event parameter is bound existentially: For [[S_{8}]]_{o} = x_{e}careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX) and in S_{11} we have everything except for the raise accented quantifier:

((S_{11}]]_{o} = λy.car′(y) ∩ λx_{i}(∃x_{e}careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX))
= λx_{i}(car′(x_{i}) ∧ (∃x_{e}careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX)))

Now, to interpret the raised quantifier, we will have to not only know the value [[S_{11}]]_{o} but also the value [[S_{11}]]_{f}. In the wellknown manner, this value is calculated like [[S_{11}]]_{o} by replacing the focussed element, here: ,,carefully“ by all possible alternatives in the context. It is sensible to take that the alternatives for an adverb of manner will again be adverbs of manner; or even that the contrary ,,non-carefully“ will be among them. So what we get is the following:

((S_{11}]]_{f} = {λx_{i}(car′(x_{i}) ∧ (∃x_{e}careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX))),

λx_{i}(car′(x_{i}) ∧ (∃x_{e}reluctant′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX))), λx_{i}(car′(x_{i}) ∧ (∃x_{e}non − careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX))), λx_{i}(car′(x_{i}) ∧ (∃x_{e}thoughtless′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX))), ...}

Now we can form the union of this set of properties and will end with the property:

λx_{i}(car′(x_{i}) ∧ (∃x_{e}(careful′(MAX, x_{e}) ∨ reluctant′(MAX, x_{e}) ∨

non − careful′(MAX, x_{e}) ∨ thoughtless′(MAX, x_{e})) ∧ wash′(x_{i}, x_{e}, MAX)))

Let us now assume that the disjunction over the possible ways of washing cars in fact exhausts all the various manners in which Max has washed cars (at least if non − careful′ is among them, this will even be a logical consequence). Then we get the property:

λx_{i}(car′(x_{i}) ∧ (∃x_{e}wash′(x_{i}, x_{e}, MAX)))

that is, we operate with the property of being a car washed by Max. Now we can interpret the top node in the tree:

((S_{13}]]_{o} = [[most]][λx_{i}(car′(x_{i}) ∧ (∃x_{e}wash′(x_{i}, x_{e}, MAX))),

λx_{i}(car′(x_{i}) ∧ (∃x_{e}careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX)))]

= MOST [λx_{i}(car′(x_{i}) ∧ (∃x_{e}wash′(x_{i}, x_{e}, MAX))),

λx_{i}(car′(x_{i}) ∧ (∃x_{e}careful′(MAX, x_{e}) ∧ wash′(x_{i}, x_{e}, MAX)))]

This will be true, iff among the cars that max washed, a mayority was washed in a careful manner - which is exactly what we predicted for the sentence.

5 Comparison with Topic-Focus Theories

I do not want to finish without passing an alternative approach to the data under discussion which makes cruical use of the notion of topic, in addition to focus. For this approach (in its various spellouts), the primary data are slightly different ones, although with the same intonation pattern as ours. Look at the following example:

5.1 Max/ kissed Amalie passionately\.

In contrast to the cases where a raise accent is placed on a quantifier, here we have a raise accented name, which evidently has the effect to evoke alternatives to that name, Max, as we would expect for ordinary focus. But for ordinary focus examples, we always get the effect that the focus, as long as it doesn’t interact with an operator that tells you otherwise, is interpreted as follows: First, I give you a set of possible alternatives to the focussed thing. Second, I am to be read as: this proposition exhausts the set of true propositions among the possible alternatives. All other alternatives are understood as being negated. So if 5.1 was to be read as a double focus example, it would say: ,,Max kissed Amalie passionately, and noone else kissed her in any other specific manner“. But this is not what we understand 5.1 to say. We read it as: ,,As for Max, we know that he kissed Amalie passionatly, and there are alternatives to Max who kiss Amalie somehow, we will tell in a minute.“ (or: ,,...which does not matter at the moment“, or: ,,...which we do not know exactly“). In a way, this reading is analoguous to the one in our examples, and in a way, it isn’t. We will compare in a minute.

Now the common observation made in work done independently by D. Buering (Buering[94]), K.v.Fintel (v.Fintel[94]) and G.Jaeger (Jaeger[94]) is the following: When looking at sentences like 5.1, there seem to be people under discussion, namely the alternatives to Max, and we discuss their behaviour where again various possibilities, indicated through the second (fall) focus are under debate. Now the fall accent for each of the alternative individuals to Max, as usual, indicates the exhaustiveness of the proposition among the things that truely can be said about Max and his mates, in the range of the focus alternatives of the utterance. So we treat this as a focus accent. The raise accent indicates that the truth or falsity of certain other propositions is still unsettled. So we call it a topic accent and evaluate it slightly different from the focus accent. In the work of v.Fintel and Jaeger, it is taken that the division of the sentence in topic and focus is a universal process that always takes place; v.Fintel’s work consists in developing a formal account for this assumption. Under this perspective the raise accent is handled as sort of a focus-within-topic and the different effects of the focus accent are accounted for by evaluating it in different surroundings. In the following, however, I want to relate to the work by Buering, because it seems to me a minimally technical, maximally transparent approach. His story goes like this: There are basically two sorts of accenting, one focus, noted as < word >_{f} , and one topic, noted as < word >_{t}. The effect of both accents is the same, namely to evoke a set of alternative values to the accented one, as in the work of Rooth (Rooth[85]). But now, the overall sentence can be evaluated with respect to focus alone - where we do not take care of the topic accent at all, and in addition the topc-focus-value of the sentence can be computed - which will be the same as if there had been double focus to be interpreted. In the example above, we can thus compute three semantic objects from the sentence:

5.2 Max_{t} kissed Amalie passionately_{f} = S

[[S]]_{o} = ˆ ∃e(KISS(MAX, AMALIE, e) ∧ P ASSIONAT E(MAX, e))
[[S]]_{f} = {ˆ ∃e(KISS(MAX, AMALIE, e) ∧ BORED(MAX, e)),
ˆ ∃e(KISS(MAX, AMALIE, e) ∧ EXCIT ED(MAX, e)),

ˆ ∃e(KISS(MAX, AMALIE, e) ∧ RELUCT ANT (MAX, e))...}

[[S]]_{tf} = {ˆ ∃e(KISS(MAX, AMALIE, e) ∧ BORED(MAX, e)),
ˆ ∃e(KISS(MAX, AMALIE, e) ∧ EXCIT ED(MAX, e)),

ˆ ∃e(KISS(MAX, AMALIE, e) ∧ RELUCT ANT (MAX, e)),

ˆ ∃e(KISS(CARL, AMALIE, e) ∧ BORED(CARL, e)),

ˆ ∃e(KISS(CARL, AMALIE, e) ∧ EXCIT ED(CARL, e)),
ˆ ∃e(KISS(CARL, AMALIE, e) ∧ RELUCT ANT (CARL, e)),

ˆ ∃e(KISS(LUDW IG, AMALIE, e) ∧ BORED(LUW IG, e)), ...}

How comes the third value, [[S]]_{tf} , into play? The rules to relate the three values say roughly: Among the values in [[S]]_{f}, what is said in [[S]]_{o} is exhaustive: it is the only true thing that can be said in this set. The value [[S]]_{tf} constitutes the global topic, what is under discussion at the moment. It can be tested against the appropriate questions, which are taken to be an ouvert manifestation of the topic. So, the appropriate topic that sentence 5.1 could be uttered in would be manifested by a question like: ,,Who kissed Amalie which way? “ or ,,How did everybody kiss Amalie? “, where a list answer will be appropriate. Formally, the value [[S]]_{tf} must be equal to the value of the topic question if the utterance is to be felicitous. The sentence then gives a partial answer to the topic question; in our example it would be: ,,As for Max, I can answer your question: he kissed Amalie passionately. And wait what about the others...“. Formally, the partiality of the answer is deduced in the following manner: First, cancel the values in [[S]]_{f} from the set [[S]]_{tf}. These propositions are out of the race, that is either stated as true, or stated as false. Then - possibly - cancel out all further propositions in [[S]]_{tf} which we deduce the truth value for, either by logical reasoning from what we already know, or possibly also by further pragmatic argumentation along gricean lines. Now the topic accenting was appropriate exactly if after all this cancelling, we do not end with the empty set, that is, if indeed part of the question under debate is still unsettled. Otherwise the ordinary focus accent that matches questions along the lines indicated in Rooth[92] would have been already sufficient. The set of propositions we end with is called the remnant topic. This is what the next utterance has to be matched against (if the interlocutors do not decide to switch the topic altogether). To give an example, in 5.2 these steps would lead us to the set:

5.3[[S]]_{tf}\[[S]]_{f} = {ˆ ∃e(KISS(LUDW IG, AMALIE, e) ∧ P ASSIONAT E(LUDW IG, e)),

ˆ ∃e(KISS(LUDW IG, AMALIE, e) ∧ EXCIT ED(LUDW IG, e)),
ˆ ∃e(KISS(LUDW IG, AMALIE, e) ∧ RELUCT ANT (LUDW IG, e)),
ˆ ∃e(KISS(CARL, AMALIE, e) ∧ P ASSIONAT E(CARL, e)),
ˆ ∃e(KISS(CARL, AMALIE, e) ∧ EXCIT ED(CARL, e)),
ˆ ∃e(KISS(CARL, AMALIE, e) ∧ RELUCT ANT (CARL, e)),
ˆ ∃e(KISS(ERNST, AMALIE, e) ∧ P ASSIONAT E(ERNST, e)), ...}

where no further reductions can be made. That is, we will predict that an appropriate next utterance could be: ,,Well, and Ernst/ kissed her excitedly\, and for the rest I do not know.“ This is intuitively the correct result.

So basically the approach distinguishes between two accents, one opening up alternatives without claiming exhaustiveness, and one giving total information within a set of alternatives. This way to proceed works remarkably well in all cases where the raise accent intuitively opens up a set of alternatives that are of the same type as the item that actually is used in the sentence. So it can also handle the longstanding examples with scope inversion, like

5.4 Everything/ doesn’t\ cause cancer.

and other cases where a quantifier is raise-accented with the meaning ,,take alternative quantifiers to stand in this place“, like in the german example (which for certain reasons cannot be reproduced in English straightforwardly)

5.5 Drei/ Aufgaben hat jeder\ Schueler geloest.

three exercises has every schoolboy solved

(Manche sogar vier, und einer hat fuenf geschafft)

= Three exercises is what every schoolboy solved (some even four, and one managed five)

What about the quantifier cases we are dealing with? One idea could be, to treat them just in the same manner as the other examples, but this idea doesn’t work out correctly. This can be seen most clearly when looking at an example with ,,half of“:

5.6 Half/ of the children wore green\ shorts.

What our semantics for this example will predict is, that it says: There are children wearing shorts, and half of them: green shorts. (The others: other colours). This is the correct meaning of the sentence; but moreover it is compatible with the other half of the children wearing, f.i., red shorts (which might even be the only alternative for green shorts) such that the sentence:

5.7 Half of the children wore red shorts.

will also be true. But this is not what the original topic approach will allow. If one takes, as suggested by Buering, the raise on ,,half of“ to evoke alternative quantifiers and start the whole business, one will predict that, as soon as ,,half of the children wore green shorts“ is recognized as exhaustively true in the set of focus alternatives of the sentence, the other propositions ,,half of the children wore red shorts“, ,,half of the children wore pink shorts“ are automatically predicted to be false. But this is not what we want. When we are told that ,,some/ of the children wore pink\ shorts“, all other ,,somes“, ,,manys“, ,,severals“ etc. are not influenced by the truth of this sentence. We know only that there is a subset in the set of shorts wearing children that consists of children wearing pink shorts. I think that there still is a unifying idea behind the examples with raise accented quantifiers (in the FNQ sense) and the other examples. The link is given by examples like:

5.8 The female/ players wore darkblue\ shorts

This parallels an example of Buering, where he suggests to immediately break down the global topic question ,,What shorts did the players wear“ from the global answers: ,,The players wore blue shorts, the players wore green shorts, the players wore red shorts, ...“ into the full list answers: ,,player 1 wore blue shorts, player 1 wore green shorts, player 2 wore blue shorts, player 3 wore red shorts, ...“ (the exact question semantic is still left unsettled in his approach). At this point one can start checking true answers, possible answers that turn out as false, and calculate the remnant topic. I in turn would like to suggest that the uniform thing happening with all such partial answers is, that the domain of individuals in question is divided by some partition. On all the partition sets, the individuals behave uniformely with repect to the property asked for (i.e. they wear the same sort of shorts). So, look at the topical question:

5.9 What sort of shorts did the players wear?

If we start with the partial answer:

5.10 The female/ players wore darkblue\ shorts.

we thereby give the beginning of a partition of the set of players into females and others. Whether the others still have to be subdivided into more partition sets is still unclear, but as for the females, we know: they wear darkblue shorts. If we start with the partial answer:

5.11 Tom/ wore green\ shorts.

we give the beginning of a partition of the set of players into singleton sets. The first such set consists of Tom alone, and he wears green shorts. Maybe it event turns out that there are bigger groups of players with uniform shorts, but as for a start, we only tell about the set consisting of Tom. The real gain of this view can be seen when we turn to the quantificational cases: Take as an answer to 5.9:

5.12 Most/ players wore white\ shorts.

What does this intuitively say, again? (- as we have asked quite often by now.) There is a subset of the set of players, not distinguished by any other feature (like being the female ones, being the set consisting of Tom, or Tom and Fred, or...) apart from being the biggest subset of players which behave uniformely with respect to the property asked for, in that they wear white shorts. This again is the beginning of a partition of the set of players, but in contrast to the other cases, it is so to speak a free partition, only dependent on shorts, not on other things. I do not intend to formalize this idea here although I think that the real truth on accents is somewhere along these lines. What we can conclude from this discussion is the following: First, also in accounts of rise- and fall-accents where a comparatively uniform interpretation of the accents is reached, the cases studied by us as FNQ still remain different. Second, nevertheless there might be integrative repairs of the shortcomings of the approach presented (which admittedly is still very in progress) which give us a global view on accents and their effects; in special these integrative solutions might be able to treat some of the points I have explicitely left open in this paper. Third, and most interestingly, one has seen now that interaction of focus with nominal quantifiers can be either treated as mainly parallel to the examples with adverbial quantifiers, or else as a phenomenon entirely based on topic and focus. When two explanations overlap on a certain range of data, one should ask whether there are other cases where the alternative approaches might compete in explanative power.

PART 2: An application in adverbial semantics

6 Adverbs of manner in German

It has often been observed that in German, predicate adverbs can turn up in various positions in the sentence (and not only in German, but I want to stay with these examples for now). Various authors have argued for both thesis, either that the standard case should be that (predicate) adverbs can be base generated in several positions, or that there should be base positions for adverbs and the varying surface orders get derived by scrambling. I would like to maintain the claim that the latter assumption makes more sense, and that certain readings for sentences with predicate adverbs have been left unnoticed up to now, due to the fact that they arise through a non-standard, namely rise-fall intonation pattern, which I have just given a semantics for.

Predicate adverbs in german can appear left of all objects, immediately following the finite verb in verb-second sentences, or immediately following the subject in verb final examples. This word order normally is relatively unproblematic in that indefinites have a normal, indefinite interpretation; and moreover are the normal word order in the terminology of Hoehle (Hoehle[82]): an accent on the appropriate NP can be the reflex of foci all the way up to the global sentence.

6.1 Maria schreibt vorsichtig einen Brief.

Maria writes carefully a letter

6.2 Ludwig hat vorsichtig alle Autos gewaschen.

Ludwig has carefully all cars washed

6.3 Hans schenkte zoegernd einer Bettlerin einen Groschen.

Hans gave  reluctantly a beggar  a dime

6.4 * Paulchen hat froehlich keine Wurst gegrillt.

Pauly  has gaily  no sausage grilled

Only one sort of examples behave pathological in this order, namely the ones which contain a negation under the adverb, like in 6.4. Conceptually it is quite clear why a sentence like 6.4 should be ill- formed. Where Paul doesn’t perform a certain action, it is senseless to predicate ,,joyfulness“ on the event that is said to not happen. Nevertheless it is not really clear how sentences like 6.4 should be ruled out. We will come back to this question later. On the other hand, predicate adverbs can appear further to the right in a sentence, as for instance in

6.5 Maria hat den Hund sorgfaeltig untersucht.

Maria has the dog  carefully  examined

6.6 ...weil Hans einen Stuhl sorgfaeltig mit roter Farbe anmalt.

(because) Hans a chair  carefully  with red paint  paints

6.7 John schmierte keine Brezel sorgfaeltig .

John buttered no Brezel carefully

6.8 ...weil Ede ein Bier hastig getrunken hat.

(because) Ede a beer hastily drunk has

In these constructions, only one thing is clear. They do not exhibit the normal word order, that is, if only one accent is to be set per sentence, this accent in the examples given will always be less than sentence focus. So, to go through one example in detail, sentence 6.8 could be mono-accented like:

6.8a Maria \ hat den Hund sorgfaeltig untersucht.

Maria has the dog  carefully  examined

6.8b Maria hat den Hund\ sorgfaeltig untersucht.

Maria has the dog  carefully  examined

6.8c Maria hat den Hund sorgfaeltig \ untersucht.

Maria has the dog  carefully  examined

6.8d Maria hat den Hund sorgfaeltig untersucht\.

Maria has the dog  carefully  examined

(I am leaving out the ,,verum-focus“ on the finite verb). The intonation 6.8a will be read as narrow focus on ,,Maria“ ( - such that Maria is contrasted with other possible examinators of the dog). 6.8b will contrast the dog with other possible objects of investigation. 6.8c will be read as stressing the manner of the investigation in question, and 6.8d will be read as reporting either what Maria did carefully to the dog, or else what Maria did with respect to the dog (where thus the accent reflects focus on ,,sorgfaeltig untersucht“). This latter intonation pattern is in the literature often given as the ,,normal“ wide focus intonation pattern for sentences with verbs of examination (compare Cinque’s example 38b. ,,Der Arzt wird den Patienten sorgfaeltig untersuchen\“, Cinque[93], p.255). I think that this is not correct. Due to several factors the feeling that no special restrictions are made to the context might arise: On one hand, the definite object has to be accomodated to the context anyway. So what is more natural than to expect that the normal course of conversation will now be: tell me what the subject (= Maria, der Doktor) did to the object (= der Hund, der Patient). This, although expectable, is nevertheless not a wide focus reading. Moreover in the doctor-patient-examples, the object ,,der Patient“ comes as expectedly with the subject ,,der Doktor“, as ,,pepper“ would come with ,,salt“, or ,,father“ with ,,son“, such that even more, these two coming together do not evoke the feeling that a non-normal division of sentence material in foreground and background is given. But take the sentences in a different context:

6.9 When we entered the sick-room, a perfect chaos expected us. The nurse was chasing a cat. The janitor was cleaning the windows, and

... der Arzt hat den Patienten sorgfaeltig untersucht.

the doctor has the patient  carefully  examined

In this context, an accent on ,,untersucht“ is not the unmarked choice, but rather a multi accent pattern like

der Arzt hat den Patienten/ sorgfaeltig/ untersucht\.

As this context breaks up the automatic association of ,,doctor“ with ,,patient“ but nevertheless asks for sentence focus (because we are listing the things that ,,happened“, when we entered the room), I take this multi-accent to be indeed the intonation pattern reflecting global focus. Analogously, the global focus intonation pattern of 6.8 would be

6.8e Maria hat den Hund/ sorgfaeltig/ untersucht\.

With this lengthy discussion of the oddities of one single example, we have already seen the additional difficulties we might come across when investigating the question: what can sentences with predicate adverbs mean in which word order (and with what sort of NPs as subject and objects). So, the factors we have to keep under control are roughly the following: (A) What is the word order of the sentence we investigate? (B) Which intonation pattern shall the sentence be read with? (C) What is the nature of the subject and object NPs? The first question normally will be easy to settle, although it will turn out to be somewhat tricky to give a coherent answer to the finer question: what is roughly the syntactic structure of the sentence? I do not want to integrate any full syntactic framework into the discussion, but only plan to give statements about the location of the VP boundary that are relevant with respect to and hopefully compatible with the mapping hypothesis of Diesing (Diesing[92]).

As for question (B), I will allow the following range of patterns: single accents which are as ,,normal“ as possible, the rise-fall patterns that come along with FNQ semantics, and the multi-accents as in the discussion of 6.8. The latter are not very well described in the literature (or I could not find the relevant references), and I take them pretheoretically to roughly work in the following way: where the structure of a sentence does not allow for one single accent to be the reflex of global focus, as an escape strategy more accents are used, and their maximal projection domains are ,,glued together“ to form a single big focus. This rough impression of course asks for further investigation. The important thing to observe is that these multi- accents, however they work, will turn out to have two effects: first, they allow for indefinite readings of indefinite NPs ^{1} which are impossible otherwise; second, they rescue sentences from sensitivities against the prediscourse that are created by narrow foci.

It will take a bit longer to make question (C) precise. I mainly want to investigate the behaviour of indefinite NPs and NPs with the determiner ,,no“ around the adverb. As for the latter, we will, apart from the two examples with ,,no“-NPs given quite commentless in the introduction, postpone this discussion until later. Indefinite NPs have been investigated comprehensively in the dissertation and book of Diesing (Diesing[92]) and I want to relate to her results.

Diesing’s theory, in a nutshell, is the following: Indefinite NPs (which subsume bare plurals for the moment) are treated as in the discourse semantic theories of Heim and Kamp (Heim[82], Kamp[81]), that is, they do not introduce an existential quantifier, but free variables which can be bound in two ways: Either by a default existential closure process where any variable that is still free gets bound existentially, or else by another quantifier that binds ,,unselectively“, mostly quantificational adverbs, but perhaps in some cases also nominal quantifiers (like in the donkey sentences of the type ,,every farmer who owns a donkey beats it“). Now, it is assumed that at LF latest, the material in the sentence is divided into parts standing in VP (possibly including the subject), and parts standing outside VP. The VP material is representend in the scope of some global quantifier, and the material outside as the restrictor. The global quantifier is either ouvertly given as a quantificational adverb, or else assumed to be implicitely present as a division into ,,what is the common background“, and ,,what is coming new“, or as a generic ,,always“-like quantifier. The striking evidence for this thesis is the observation that, in German, it is indeed the case that indefinites an adverb like ,,usually“ quantifies over are always found to the left of the adverb, and indefinite NPs that only come in the scope of the quantifier always stand to the right of it. If the adverbial Quantifier is taken to mark the VP boundary, then the mapping hypothesis will predict exactly this behaviour. To see a minimal paradigma, take

6.10 ... weil immer ein Clown eine Torte fallenlaesst.

(because) always a clown a creamcake drops

6.11 ... weil ein Clown immer eine Torte fallenlaesst.

(because) a clown always a creamcake drops

6.12 ... weil ein Clown eine Torte immer fallenlaesst.

(because) a clown a creamcake always drops

1I will use the term ,,indefinite“ in the following way: An indefinte NP is one of the form ,,a <noun>“, or in German ,,ein(e) <noun>“. An indefinite NP can be read partitively, or generically, or indefinitely. So, it is not tautological to say that the indefinite NP was read indefinite.

6.10 says that always (in appropriate situations, maybe in a circus show), there is a clown who drops a creamcake. So it is representend ^{2} as

6.10′ALW AY Ss(appropriate(s), ∃x∃y(CLOW N(x, s) ∧ CAKE(y, s) ∧ DROP (x, y, s)))

6.11 reports in turn that what a Clown always does, in an appropriate situation, is: dropping a creamcake. This is reflected in

6.11′ALW AY Sx, s(appropriate(s) ∧ CLOW N(x, s), ∃y(CAKE(y, s) ∧ DROP (x, y, s)))

6.12 finally tells that what a clown always does to a creamcake is, he drops it. So we have

6.12′ALW AY Sx, y, s(appropriate(s) ∧ CLOW N(x, s) ∧ CAKE(y, s),

DROP (x, y, s))

Moreover if an indefinite NP goes out of VP, which also can be signalled by other means, like particles, and thus evades existential closure that would normally have bound the free variable introduced in the interpretation of the indefinite NP, something else has to happen. Either then the indefinite is read generically, so that at the VP boundary something like a silent ,,always“ could be understood. Or else it is read what Diesing calls ,,partitive“ and what I want to take as being reflected by a rise fall intonation pattern on the indefinite article + somewhere else and interpret as an FNQ reading. In any case, the effect is that the indefinite NP is not read as a mere existential statement but has its own restrictor and scope, either to be used in FNQ interpretation or noted as an indication for this NP being a (strong) quantificational one. An example for this thing to happen is:

6.13 ...daß ein Kind ja weint.

(that) a/one child PRT cries

6.14 ... daß ein Kind ja weint\. 6.15 ... daß ein/ Kind ja weint\.

The particle ,,ja“ has been shown to mark the VP border in German. And indeed, sentence 6.13 can not be read as to say that there was a child, and that this child cried, but only that either in general, children cry, or else that one of a certain crowd of children cried. This is what Diesing observed and covered with her theory. We now are in the lucky state that, at least for the partitive reading, we can give a more concrete explanation of what is going on. The ,,ein“ is interpreted as quantificational ,,ein“, and the rise fall pattern is interpreted as focus (on ,,weint“) interacting with this nominal quantifier, as given in the preceding chapters. It is not equally clear what is going on in the generic reading that comes along with the intonation pattern as I gave in 6.14. Nevertheless we are able to make the partitive reading prominent by shifting intonation in all the examples in question, and moreover it turns out that without the raise on the ,,ein“ at least, no partitivity can be understood (see also Jaeger[94]).

Let us against this background investigate the question whether predicate adverbs evoke comparative effects, and whether the mapping hypothesis of Diesing can tell anything in these cases. When looking at the cruical examples, it immediately turns out that two things hold. On the one hand, predicate adverbs like ,,froehlich “ (joyfully), ,,hastig“ (hastily), ,,vorsichtig“ (carefully), ... cannot be the VP boundary because indefinite subject NPs can well be read indefinite left to them:

2I give the predicates CLOW N and CAKE as binary, with an argument for the situation parameter. They are to be read as: x is a clown or cake in s. This is necessary because of the partiality inherent in situation semantics. I refer the reader to Muskens[91] for a suggestion how a simple situation semantics could work.

6.16 ... weil ein Busfahrer vorsichtig einen Verletzten

(because) a bus driver carefully an injured

aus dem Unfallauto gezogen hat.

out of the car drawn has

6.17 ...weil ein Busfahrer den Verletzten vorsichtig

(because) a bus driver  the injured  carefully

aus dem Auto gezogen hat.

out of the car drawn has.

The NP ,,ein Busfahrer“ can perfectly be read indefinite, as ,,some bus driver“ who neither is the typical, nor one out of a set of known bus drivers. This contrasts with the behaviour of ,,ein Clown“ in the sentence adverbial example 6.10. It is not surprising because the subject NP can never be located right to the predicate adverb (PA in the following), so that if ever an indefinite subject is to go with a PA at all, it always must be left to it ^{3}. But for indefinite object NPs the picture looks quite different. They undergo the typical shifts in meaning whenever they come left of the PA. So, in the sentence

6.18 ...weil Franz ein Bier vorsichtig getrunken hat.

(because) Franz a beer carefully drunk has

the object ,,ein Bier“ can at first hand be understood either as generic, such that the sentence says: ,,What Franz always did with a beer is: drink it carefully.“, or as partitive in the intonation

6.18a ... weil Franz ein/ Bier vorsichtig\ getrunken hat (vorsichtig getrunken\ hat).

telling that one of the beers he consumed, Franz drank carefully. (Alternatively: one of a set of beers known, Franz drank carefully.) So, the first sort of examples seem to indicate that the subject left to the PA is within the VP (in the light of Diesings hypothesis), whereas the second sort of examples would show that the object NP between the subject and the PA is outside the VP. Now, 6.18 is chosen with a proper name as subject, so that we do not see what an indefinite subject would do in that case. One might hypothesize that maybe the object was not ,,squeezed in“ between subject and PA, but that instead it was adjoined outside VP and the subject to its left. Thus, 6.18 should have a structure like:

6.18b ... weil [_{s}Franz_{i}[_{s}ein Bier_{j}[_{V P} t_{i}[vorsichtigt_{j}...getrunken hat]]]]

Under that assumption one would expect that indefinite subjects in such examples underwent the same shifts in meaning. Now, look at 6.19 ... weil ein Busfahrer einen Verletzten vorsichtig aus dem Auto

3

? ... weil vorsichtig ein Busfahrer einen Verletzten aus dem Auto gezogen hat.

can only be read with the adverb in narrow focus (topicalized?)

(because) a bus driver  an injured  carefully out of the car

gezogen hat.

drawn has

To make things clear, I give one partitive intonation pattern for the object NP, and one generic. So we have

6.19a ... weil ein Busfahrer einen/ Verletzten [ vorsichtig aus dem Auto

\ gezogen hat ]_{f}. 6.19b ... weil ein Busfahrer einen Verletzten/ vorsichtig aus dem Auto

gezogen hat.

6.19a is the partitive pattern, 6.19b the generic - where further accents are optional (and evoke respective focal shifts). Now the thing is: In 6.19b indeed nothing rescues the bus driver from being understood either partitive or also generic. So, 6.19b is either to be read as

,,One of the bus drivers always when presented with an injured (lying in a car, probably): drew him carefully out of the car.“

or as doubly generic:

,,whenever a bus driver hits on an injured (in a car), what he does is: drawing him out of the car, carefully.“

The properties ascribed to the bus driver may vary with changes in intonation in the VP, but the bus driver himself cannot be indefinite any more. But in 6.19a on the other hand, it is still possible to have the bus driver indefinite and understand:

,,So, there was a bus driver who drew one of the injured out of the car.“

Thus according to the mapping hypothesis, the NP ,,ein Busfahrer“ cannot stand outside VP . Now, this is not totally incompatible with what we find elsewhere, or in Diesing. We can claim that there are two ways to get an object to the left of the PA: If the object is to be read generically, it must stand left to some silent GEN operator at the VP boundary, which I want to assume to have a semantics like a liberal ALW AY S quantifier, to keep things easy. Thus 6.19b has a structure like:

6.20...weil[_{s}ein Busfahrer_{i}[_{s}einen Verletzten_{j}[_{s}GEN[t_{i}[_{V P} vorsichtigt_{j}

aus dem Auto gezogen hat]]]]

with stressed and thus quantificational ,,ein“ in the subject, or generic subject and object NP. So this will explain why the subject NP in 6.19b is barred from getting an indefinite interpretation, but is fine either when also bound by the silent GEN, or quantifying on its own. If the object is to be read quantificationally on the other hand, we may claim that in this case it is indeed squeezed in between subject and PA and so still within VP - but not in its base position. The respective structure will look like this:

6.21...weil[_{V P} ein Busfahrer[_{V P} einen Verletzten_{i}[vorsichtigt_{i}

aus dem Auto gezogen hat]]]

This is fine apart from leaving us the duty to tell something about, why the object NP - still being comfortably within VP - is barred from an indefinite reading then. But well - it is not, really. As we have seen already in the introductory examples in this section, there is the multy accent pattern that seems to heal everything. Let us try in our example:

6.22 ... weil ein Busfahrer einen Verletzten/ vorsichtig/ aus dem Auto\ gezogen hat.

So what: indeed, now again we can understand this as saying:

,,Well, there was this horrible accident. Now there was also a bus driver standing on the road, and what he did was, he carefully drew an injured out of the car. (I couldn’t even see whether there were more people in the car, injured or not).“

Also when the object NP receives an accent alone, which is interpreted as narrow focus, it is possible to read it indefinitely. In this case, however, it is better to have a definite subject NP, because it would otherwise be an indefinite standing in the background part of the sentence. This evokes further difficulties, as we will see in a moment. So the repair is like:

6.23 ... weil Franz einen Verletzten\ vorsichtig aus dem Auto gezogen hat..

This means something like (by interpreting the narrow focus as contrastive)

,,It was not a corpse or a intact person, no, Franz carefully drew an injured person out of the car.“

Again, the indefinite NP ,,einen Verletzten“ is indeed interpreted indefinitely - although it stands in contrast to other possible candidates to be drewn out of the car, it still means: There was this injured person, and Franz rescued him out of the car.

To summarize what we found up to now: Subject NPs can appear in indefinite readings to the left of a predicate adverb. Thus PAs cannot automatically be taken to mark the VP boundary in German. Object NPs at first hand seem to be read quantificationally or generically when they appear left to a PA. This suggests that their position is outside VP, in the light of Diesing’s theory. A generic object NP left to a PA forces the subject NP to be read generically or partitively, too. So here it is consistent to assume an underlying structure like in 6.20, where both subject and object NP are moved out of the VP. A partitive object NP left to a PA however does not affect the indefinite reading of the subject. We thus assumed that alternatively to the structure given in 6.20 also the pattern in 6.21 would be possible. What about reading the object NP indefinitely in these cases? At this point we hypothesized about an entirely new factor that is in play. What if it was necessary for an indefinite NP to be read indefinitely that it was somewhere in the focus part of the sentence? With the non-normal word order that we diagnosed for sentences where the PA does not appear immediately after the subject, it would be clear that without further effort the object NP was not in the focus part of the sentence. If we make this further intonatoric efforts however, by spreading multiaccents over the sentence or narrow focus accent on the indefinite itself, we are again allowed to interpret the indefinite NP as indefinite.

This would be a nice claim to test agains the data, but unfortunately a further difficulty arises through examples with FNQ. There, it seems, our claim that indefinite NPs must always be in the focussed part of a sentence fails. The example is like this:

6.24 Die meisten/ Bauern haben andaechtig\ einen Schnaps getrunken.

the most  farmers have devoutly  a schnaps  drunk

6.25 Ein/ Bauer hat andaechtig\ einen Schnaps getrunken.

one farmer has devoutly  a schnaps  drunk

Here we have the wellknown raise accent on the determiners indicating an FNQ construction. The fall lies on the adverb and can only reflect narrow focus. Thus, the indefinite object NP (well stored in VP) is not in focus. It should thus be read somehow non-indefinite. But nevertheless the sentences say something like:

6.24a Most of the farmers who were involved in drinking a schnaps did it devoutly. 6.25a One of the farmers who were involved in drinking a schnaps did it devoutly.

A more elaborate paraphrase which is closer to the final logical representation runs like:

6.24b For most of the farmers for which there was a schnaps they drank there was a schnaps they drank devoutly. 6.25b For one of the farmers for which there was a schnaps they drank there was a schnaps that he drank devoutly.

Thus we see that the indefinite ,,einen Schnaps“ is treated like anything else standing in the focus-background part of the sentence in FNQ-constructions: It basically goes into the restrictor. The feeling that the schnaps to be drunk was the same in restrictor and scope is the same as in the examples with Max washing cars that we in depht discussed in section?. What we called the one-one-reding there with one washing by Max per car, here turns up as the one-one-relation between schnapses and farmers who drank it, probably mediated through the one-one-relation between farmers and drinkings (where only one drinking has occured, only one schnaps can be drunk per farmer). But what should interest us now is the fact that the indefinite object in its indefinite reading seems to be quite fine outside the focus! Why didn’t we come across data like these in the above examples? Let us change the word order in 6.24 and 6.25 and put the object in front of the PA:

6.26 Die meisten/ Bauern haben einen Schnaps andaechtig \ getrunken.

the most  farmers have a schnaps  devoutly  drunk

6.27 Ein/ Bauer hat einen Schnaps andaechtig \ getrunken.

one farmer has a schnaps  devoutly  drunk

It turns out that the readings we find here are in line with the previous hypothesis that indefinite NPs cannot be interpreted indefinitely when they are in the background part of a sentence. The paraphrases of (26) and (27) are:

6.26a For most of the farmers who’s regularities in schnaps- drinking behaviour are under debate the typical schnaps drinking went on in a devout manner. 6.27a For one of the farmers who’s regularities in schnaps- drinking behaviour are under debate the typical schnaps drinking went on in a devout manner.

So, however the details of the quantification might look like, in any case the schnapses are read generically ^{4}. And we can even explain why that is: The indefinite NP ,,einen Schnaps“ is not in a position where it can be in focus on behalve of another accent - so is not in focus. It does not carry an accent on the determiner (no intonational variatiuon is assumed between the raise and the fall) - so cannot be read partitive. So the only interpretation left is the generic - which indeed is understood. What I want to suggest as an explanation for why 6.26 and 6.27 work as they do is the following: The indefinite object ,,einen Schnaps“ is somehow incorporated into the verb meaning to form a complex predicate ,,einen Schnaps trinken“ (drink a schnaps). Other examples of this kind are pairs of intransitive verbs and transitive verbs with a certain object which say the same, such that the transitive verb plus object is synonymous to the intransitive one. Here the (indefinite) objects never have an independent status. In German we have

4These are by the way the only sort of eamples where unselective binding by the nominal quantifier might be suspected. Several observations stand against this claim, however. First in (6.27) this would amount to saying that one farmer drank one schnaps, devoutly - which is certainly false. It is more correct and elegant a solution to diagnose GEN at the VP boundary. Second, as indefinites not only by word order, but also by focus-backgrounding stand in restrictors of the nominal quantifiers, it would have to be explained why in (6.24)/(6.25) the respective ,,unselective“ quantifications do not take place.

eine Verbeugung machen = sich verbeugen (to bow)

ein Aquarell malen = aquarellieren (to draw a water painting) einen Kopfstand machen = kopfstehen (to stand on your head)

Now, one has to say what it means ,,to be incorporated“. Of course in the course of semantic interpretation - as long as the framework is compositional - we can always compute a semantic object from the meaning of the verb plus the object, and then further combin it with other items in the sentence. Thus incorporation semanticaly must be more than just semantic combination. One difference between semantic incorporation and semantic combination might be, for instance, that the conditions normally to be checked with indefinites are lost, because the indefinite does no longer appear as an autonomous item to be interpreted, but has been melted into the meaning of the complex predicate. So, where no indefinite is interpreted, we no longer need to check whether it is in focus or not. I am well aware that this is a hypothesis that has to be tested against other things one expects from incorporations. One urgent question to be answered is: What are the syntactic restrictions to this incorporation process?

As we saw, it was possible to incorporate in 6.24 and 6.25, whereas not in 6.26 and 6.27. Is this due to the fact that in the former ecxamples, the object NP was in its base position, and not in the latter examples? Alternatively the cruical factor might be: in one case there is a syntactic constituent dominating exactly the ingredients of the complex predicate but not in the other case. Unfortunately I must leave this question open at the time being. The overall hypothesis now stands as:

Indefinite NPs are good indefinite only as long as they are within VP and in focus, except if their indefiniteness can be rescued by incorporating them into a complex predicate, where they may be in focus or not.

Stated in this form, it becomes evident that as long as no independend criteria are worked out for when we may incorporate and when not, this gives us the zero theory. This will be one urgent point of investigation, and any evidence for or against this hypothesis will be valuable.

The choice and order of examples discussed in this section was mainly driven by the plot of the semantic arguments. The reader may be left with a feeling that ,,anything goes“. To give the data structured in a different way, I have added in the appendix a systematic overview over the things that can happen with

• sentences in normal word order,

• with object NP squeezed in between subject NP and PA,

• with object NP set left to both, subject and PA (which I have not mentioned up to now) and

• with both, object and subject adjoined left to the VP boundary.

Although I was not able to give everything that these sentences can report I tried to be complete at least in readings that would matter in the above discussions.

7 Adverbs and Negation

What I want to discuss in this section is mainly the minimal pair

7.1 * John hat vorsichtig keine Brezel geschmiert.

John has carefully no brezel buttered

7.2 John hat keine Brezel vorsichtig geschmiert.

John has carefully no brezel buttered

or more generally: how can the unacceptability of predicate adverbs in certain negated contexts be explained? A first explanation in the light of 7.1 and 7.2 could be, to claim that in German, there was a fixed position for negation, a NegP, as has been argued by Buering (Buering[93]), and that all negative determiners are agglomerated negation plus indefinite article. The predicate adverbs would then have a position under this negation phrase and thus, 7.1 would be bad already for syntactic reasons. However, if we look at 7.2 carefully, we note that what indeed is in play here is an FNQ construction with a raise understood (maybe not so prominently pronounced though) on ,,keine“ and a fall on ,,vorsichtig“ . I think this effect can be reproduced in English also, with the famous example ,,John sliced no bagel carefully“ . So what 7.2 really is understood like is: ,,With none of the Brezeln John buttered, he did the buttering carefully“ . Conceptually this makes a big difference. Whereas in 7.1 it is reported that something did not happen, namely Brezel butterings by John, and it is thus quite senseless to modify the manner in which what has not happened has not happened - in 7.2 we really talk about a certain set of events, namely the Brezel butterings done by John. For this restricted set of events, it makes well sense to look at each of them and predicate non-carefullness of it. The syntactic explanation for the difference between 7.1 and 7.2 does not integrate this conceptual difference. As for why the FNQ reading would be preferred it could only say that, the direct logical representation of 7.2 will be like

NOx(BREZEL(x), ∃e(BUT T ER(JOHN, x, e) ∧ CAREF UL(JOHN, e)))

This in itself is well compatible with there no Brezeln being buttered by John at all. So could it be only due to world knowledge reasoning that we nevertheless understand Brezeln as being buttered? Under this assumption, the FNQ reading would only make explicit what we understand the sentence to say anyway. But I want to show that this is not the case, and that at least for German the meaning of 7.2 reported in the literature up to now has always been the reading of the FNQ construction - because this is the only intonation pattern that is comparatively ,,normal“ and interpretable at all! How can we test this? Remember that ,,no“ /,,kein“ like ,,a“ /,,ein“ is also a weak determiner and so might fall under the intonatoric processes discussed above. So the first thing we know is, that there is no normal accent and that we always have to make intonatoric efforts to save the indefinite (weak) reading of the NP. If the meaning of sentence 7.2 above and especially the fact that Brezeln are understood as buttered, was due to world knowledge reasoning, and 7.1 out only for syntactic reasons, then it cannot affect the acceptability or interpretation of 7.2 to enforce the indefinite reading of ,,keine Brezel“ . If however the fact that it is a certain set of events we talk about is cruical for the felicitous application of the adverb, then things might change when the FNQ reading is barred by multi-accents. And indeed, when we try and multiaccent sentence 7.2, it becomes unacceptable again:

7.3 *John hat keine Brezel/ vorsichtig/ geschmiert\.

Thus where in other cases with indefinites, this accent pattern restored indefiniteness for indefinite NPs, in the negation case, it restores inacceptability for inacceptable cases. It cannot have been just because we know the world and what a sensible statement would look like, that in 7.2 at first hand we read there being butterings going on and thus understand what would also be the FNQ reading. Now example 7.3 has exactly the same word order as the correct sentence 7.2. Even more, as we have argued above for indefinites, it makes sense to assume that the underlying syntactic structure of both examples is roughly identical. So a syntactic explanation for the oddity of 7.1 could only say something about 7.3 like: maybe the sentences 7.2 / 7.3 are indeed derived from the incorrect word order in 7.1, but only 7.3, not 7.2, is also interpreted in this incorrect word order - such that 7.3 will be out. 7.2 is rescued somehow, possibly then because if it were interpreted in the original position, the determiner to interact with focus would no longer c-command the focussed item and thus could not interact with focus (as is easy to see in the mechanism given above). This explanation might work, but one is uncertain again why one should assume something like this - except for it rules out the correct cases. Conceptually however, the case is somewhat clearer (although not entirely). When presented with 7.3 we are told what John has done to no Brezel. Namely buttered carefully. Now, is this because he has buttered none? Then it will be at least under normal circumstances odd to say that this has happened carefully. Or is it because he has buttered all smeary? Then the information structure would not match the things that have happened. The FNQ reading would again be more appropriate. So as we have seen, the only thing that might rescue the syntactic approach might be: well, maybe the multi-accent pattern as in 7.3 and all the cases above does not only tell us something about the division of sentence material into focus and background of the sentence, but even forces us to interpret the sentence as if the normal word order would stand there. For the indefinite objects this makes them safely indefinite, and for the uninterpretable adverb under negation, this might again give us the unallowed constellation of adverb-over- negation. So let us investigate what the effects of multi-accents are. What we are looking for are other cases where we can place multi- accents on sentences like the ones above. Then it will have to be checked whether the object NP has scope over the adverb or vice versa, when multi-accents are interpreted. The problem is that in most cases predicate adverbs do not show scope, or at least not in a clearly traceable way. But what we can test are local adverbials like ,,ueberall“ (everywhere), and opacity inducing adverbs like ,,absichtlich“ /,,unabsichtlich“ (willingly/unwillingly). Let me start with the local adverbs. For local adverbials the tests for indefinites and intonation can be made in the same way as with the PAs in the last section, and they show the same behaviour. So in special object NPs show analogous behaviour to the left and to the right of local adverbials as with PAs. Now consider:

7.4 ... weil Otto ueberall in Europa eine Spionin verfolgt hat.

(because) Otto everywhere in Europe a spy persecuted has

7.5 ... weil Otto eine Spionin ueberall in Europa verfolgt hat.

(because) Otto a spy everywhere in Europe persecuted has

(4) in its most normal intonation (accent on ,,Spionin“ ) tells us what Otto did in all places: prosecute a spy. It is fine for this sentence to assume that in each european country it was a different spy that Otto chased. In 7.5 the first thing to note is that the usual intonatoric complications arise. To disambiguate let us look at one partitive reading example, at a generic pattern, and the multi-accent pattern:

7.5a ... weil Otto eine/ Spionin ueberall\ in Europa verfolgt hat.

7.5b ... weil Otto eine Spionin/ ueberall in Europa verfolgt\ hat. 7.5c ... weil Otto eine Spionin/ ueberall/ in Europa verfolgt\ hat.

Intonation 7.5a gives a partitive reading that says roughly: ,,One of the spys chased by Otto was persecuted even all over Europe“ . Intonation 7.5b makes the generic reading possible and says: ,,What Otto did always when he came over a spy was: prosecute her all over Europe.“ For 7.5c we note again that neither we talk about the typical female spy Otto meets or gets knowledge of, nor does the sentence imply in any way that there are more female spys in the world. However in this case it is not possible to assume that Otto prosecutes several female spys. 7.5c tells us that there was this one female spy that Otto prosecuted - all over Europe. This clearly is not the reading in the base order of the sentence like in 7.4. So we find that: multi-accents do not force us to interpret sentences in the base word order, but only have effects on the focus-background division of the sentence. The data with opacity inducing adverbs confirm this observation.Take examples like:

7.6 Paula hat absichtlich einen Ostfriesen geheiratet

Paula has willingly  an East Frisian  married

7.7 Paula hat unabsichtlich einen Ostfriesen geheiratet

Paula has unwillingly  an East Frisian  married

I do not intend to give a full semantic treatment of ,,absichtlich“ /,,unabsichtlich“ which would be an interesting matter in itself, as probably both, opacity and focus phenomena have to be accounted for. My main concern is the question if the provenience of the husband of Paula can be subject to her intensions or not. In the above examples, it clearly can: In 7.6 it is possible to understand that Paula always had the plan to marry someone who came from East Frisia (they are said to be quiet people, which Paula finds convenient). In 7.7 where Paula unwillingly married an Ostfriesian, she might be surprised afterwards when she is told that she has just married now - but she might also be surprised because she is told that the man who she in full consciousness has just married stems from East Frisia. On the other hand, when word order is changed, the plans and surprises of Paula change also.

7.8 Paula hat einen Ostfriesen absichtlich geheiratet

Paula has an East Frisian  willingly  married

7.9 Paula hat einen Ostfriesen unabsichtlich geheiratet

Paula has an East Frisian  unwillingly  married

Again we have to distinguish three possibilities to read the indefinite object. The generic reading does not make much sense here, because it will say something about Paula habitually marrying eastfrisian men - be it willingly or unwillingly. The partitive reading with fall on ,,absichtlich“ /,,unabsichtlich“ will be understood as: One of the Ostfresians that Paula married, she married willingly/unwillingly. This opens the somewhat strange picture that Paula has a harem of Ostfresians all of which she has married on purpose except for one (or resp.: all of which she married by accident except for one). Nevertheless: Subject to her will or surprise can no longer be the origin of the husbands, but only the fact of the marriage. Finally, when multi accents are placed on the sentence like:

7.10 Paula hat einen Ostfriesen/ unabsichtlich/ geheiratet\. 7.11 Paula hat einen Ostfriesen/ absichtlich/ geheiratet\.

Paula no longer lives in polygamy. There is this one East Frisian, who she has married, no other East Frisians are present and no other marriages have taken place. But however, in 7.10 the unwillingness of Paula can only hold with respect to the marriage. So she may well have known that this guy here is from East Frisia but is thunderstruck when being told that she has just married him. 7.11 on the other hand is compatible with Paula not knowing at all that her new husband comes from East Frisia, so this has not been part of her plans. The marriage itself however is always the subject of Paulas intensions. So even though the NP ,,einen Ostfriesen“ is read indefinite due to the accenting, it is nevertheless out of the scope of ,,absichtlich“ /,,unabsichtlich“ .

These two constructions suggest that placing multi-accents on sentences indeed only affects the focal structure of the sentence but has no effect on the order of interpretation of the items. The indefinite objects in both cases did not get back in the scope of the adverbials as they were in their base positions, but kept having wide scope in spite of their regained indefiniteness.

So we are left uncertain as to why the differences in acceptability of 7.1 - 7.3 arise. We have seen that a mere syntactic explanation does not apply - or at least not that easily. So it makes sense to look for a way to set down the conceptual intuitions about these sentences in a formal manner. I am embarrassed to admit that I have no further workedout answers at hand at this point.

8 Appendix

I give an overview over sentences with predicate adverbs in various word orders with the position of the VP boundary as discussed in section 6.

A. Base order

...[_{V} P NP_{subj}[_{V P} PA[_{V P} NP_{obj}... V ]]]

The examples:

8.1 ... weil ein Mann gierig ein Bier\ getrunken hat.

(because) a man  greedily a beer  drunk  has

,,There was a man who greedily drank a beer“ - sentence in normal word order with normal accent, both, subject and object NP indefinite.

Base order and FNQ construction with subject determiner and some other focus; that is: subject NP partitive, object NP (mostly) indefinite:

8.2 ... weil ein/ Mann [gierig\]_{f} ein Bier getrunken hat.

,,One of the men drinking a beer did it greedily“ . Object NP indefinite, although in the background part of the sentence, compare example 6.24, 6.25 in section 6.

8.3 ... weil ein/ Mann gierig [ein Bier\]_{f} getrunken h̄at.

,,One of the men who greedily drank something greedily drank a beer“ . Object NP indefinite and in focus.

8.4 ... weil ein/ Mann gierig [ ein Bier\ ḡetrunken h̄at ]_{f}

,,One of the men who did something greedily did: greedily drink a beer“ . Object NP indefinite and in focus.

8.5 ... weil ein/ Mann [ gierig ein Bier\ getrunken hat ]_{f}

,,One of the men in question did: greedily drink a beer“ . Object NP indefinite and in focus. It is consistent with the overall picture, although I never mentioned it above, to assume that even in these cases of partitivity FNQ is at work - although the disjunction of alternative properties of the men in question can only give somewhat vague: have some property or other that is relevant in the context. Still this can be seen as a paraphrase of ,,being one of the set of men under discussion at the moment“ .

8.6 ?*... weil ein/ Mann gierig ein Bier getrunken\ hat.

This constellation is not interpretable. It seems to be one of the cases where the above postulated incorporation process to rescue the indefinite object as indefinite would fail. A generic interpretation of ,,ein Bier“ is out because the object NP visibly is inside VP. A partitive rescue however is fine with the shifted intonation pattern like

8.7 ... weil ein/ Mann gierig ein/ Bier getrunken\ hat.

,,One of the men greedily drank one of the beers“ . - Two partitive indefinites. Do not ask me for a treatment of double raise accent.

B. Subject out of VP

...[_{S} NP_{i,subj} [_{V P} GEN [_{V P} t_{i} [_{V P} PA [_{V P} NP_{obj}...V ]]]

The visible word order is the same as under A, but in accordance to the claim that generic material is outside VP, I give this rought structure for generic subjects.

8.8 ... weil ein Mann gierig ein Bier trinkt.

,,What a typical man always does (in appropriate, or less appropriate situations) is: greedily drink a beer“ . - generic subject NP, indefinite object NP. I refrain from giving accents here. I am uncertain about how to intonate generic indefinite NPs. In some cases accenting vs. deaccenting has been observed to mark the borderline between generic and indefinite reading, like in ,,Ein Wildschwein\ lebt im Wald“ (= ,,There lives a boar in the forest“ ) compared with ,,Ein Wildschwein lebt im Wald\“ (= ,,A typical boar always lives in the forest“ ). Evidently it is too simple, though, to assume that genericity and deaccentedness are equivalent for indefinites. So, in 8.8 ,,Mann“ clearly will get an accent in a normal reading of the sentence; if noted with a slash however, it strikes me so prominent that I am seduced to read the typical man as contrasted to the typical dog, or woman, or whatever. This is why I leave out accents in these cases - try to get a native speaker of German who can read aloud, if you need them...

C. Object within VP between adverb and subject NP

...[_{V P} NP_{subj} [_{V P} NP_{j,obj} [_{V P} PA [_{V P} t_{j}...V ]]]

8.9 ... weil ein Mann ein Bier/ gierig/ trinkt\

(because) a man  a beer  greedily drinks

,,There is a man who greedily drinks a beer“ - both, subject and object NP are read indefinite. This is what we had as multi-accent-pattern in the discussions in section 9.

8.10 ... weil ein Mann ein/ Bier gierig trinkt\f.

,,For one of the beers, there is some man who greedily drinks it“ - subject NP indefinite, object NP read partitive. The accent on the verb reflects a wide focus, as indicated. It seems to be worse, for reasons still unclear to me, to put a narrow focus either on ,,trinkt“ or on ,,gierig“ , as in

8.11 ?... weil ein Mann ein/ Bier gierig\f trinkt.

and still expect the subject NP to be indefinite. This is unexpected because we would assume the FNQ business going on without even touching the indefinite subject NP, which comes in only later.

8.12 ?*... weil ein/ Mann ein Bier/ gierig/ trinkt\f.

Although in neutral sentences the multi-accent pattern works as if there was wide focus again, this seems not possible in interaction with focus sensitive nominal quantifiers. So this sentence cannot, as it should, mean: ,,One of the men greedily drank a beer“ . It is either understood as the typical behaviour of one of the men against (generic) beer (!note again that the generic indefinite is accented then), namely: he greedily drinks any beer he comes across. Or it may be understood as an unclean realization of the double raise as in

8.13 ... weil ein/ Mann ein/ Bier gierig trinkt\.

,,One of the men drank one of the beers, greedily“ . Finally note the absence of a possible reading with generic object and indefinite subject. One might speculate that this is not only out by technical reasons, but even reflects the deeper insight that showing a regular behaviour against f.i. the typical beer, might count as stable a property as ,,being intelligent“ and thus even for deeper reasons is impossible to predicate over an indefinite, in the same way as other individual level predicates are.

D. Subject and object both out of VP

...[_{S} NP_{i,subj} [_{S} NP_{j,obj} [_{V P} GEN [_{V P} t_{i} [_{V P} PA [_{V P} t_{j}... V ]]]

The examples are the following:

8.14 ... weil ein Mann ein Bier gierig trinkt.

,,What a typical man does to a typical beer is: greedily drink it.“ Both NPs are understood generically. It is not entirely clear to me whether the silent GEN operator interacts with focus in the same manner as an ouvert ,,always“ would do. So possibly by setting narrow focus on some item in the VP one can even further influence the computation of restrictor and scope of the generic operator. But this is a different phenonenon.

8.15 ... weil ein/ Mann ein Bier [ gierig\]_{f} trinkt.

,,One of the men who had certain beer drinking habits always drunk the typical beer greedily.“ This is the case we had in examples 6.20, 6.24a and 6.25a. Now that the indefinite object NP is outside VP, it can be bound by GEN and receive its generic interpretation, whereas the subject is understood partitive.

8.16... weil ein/ Mann ein Bier [ gierig trinkt\]_{f}.

,,One of the men who show certain habits towards typical beer does trink the typical beer when presented with it.“ This is the same structure as the example above, only with a wider focus.

8.17 ... weil ein/ Mann ein Bier gierig [ trinkt\]_{f}.

,,One of the men who did something greedy to a typical beet did: greedily drink the typical beer he is presented with.“ - This example is somewhat odd because it is hard to imagine a context where the question is: what do men do greedily to typical beers? Nevertheless the oddity seems to me to stem from our world knowledge, not from the oddity of the construction as such. A better example might be:

8.17a (What did nurses in the first world war do with Frantzbranntwein to injured people? - Well I know that...)

... dass eine/ Krankenschwester einen Verletzten mit Frantzbranntwein

(that) one nurse  an injured  with Frantzbranntwein

uebergossen\ hat.

poured has

,,One of the nurses always when treating an injured with Frantzbranntwein, in fact poured it over the guy“ . A very tricky question arising here is: what does the fall interact with? - with the GEN-quantifier or with the nominal quantifier, or with both - and how? We will not pursue this here.

8.18 ... weil ein Mann ein/ Bier gierig\ trinkt.

,,The typical man always drinks one of the beers greedily“ - generic subject and partitive object NP. The sentence suggests that it is always a certain beer or even beer brand that is greedily swallowed by the typical man - the first beer of the evening, or the brand ,,Lausitzer Dunkel“ . I have nothing to say on that. 8.18 could also arise in the underlying structure with subject NP out of VP, but object NP still within, like in C. Nothing else however indicates that such structures arise. In special it is impossible to have the given word order , generic subject, and indefinite object, no matter how accented. So there is no reading of

8.19 ... weil ein Mann ein Bier gierig getrunken hat

that is like: ,,What the typical man always does is: greedily drink a beer“ , like in (8).

E. Object left of subject, VP boundary in between.

...[_{S} NP_{j,obj} [_{V P} GEN [_{V P} NP_{subj} [_{V P} PA [_{V P} t_{j}... V ]]]

8.20 ...weil einen Alpengipfel ein Adler scharfsichtig überwacht.

(because) a mountainpeak an eagle sharp-eyedly surveys

,,The typical mountain peak in the Alpes is always surveyed sharp- eyedly by an eagle.“ - generic object, indefinite subject NP. It is somewhat hard to find examples that allow for such a reading. For instance with our long used example of men drinking beers, it is somewhat difficult to understand, in the analogous construction, that what always happens to a typical beer is, getting swallowed greedily by a man. Although cetrainly plausible, it is not the best way to express this observation by ,,...weil ein Bier ein Mann gierig trinkt“ . Maybe the verb type (as in the classification in Diesing [92]) plays a role, too.

8.21 ...weil einen Verkehrssuender ein/ Polizist sadistisch\ bestraft.

(because) a traffic offender one policeman sadistically punishes

,,For the typical traffic offender, there will always be one of the policemen who, when punishing him, does it sadistically“ - generic object NP, partitive subject NP. Again, it is suggested that it is one certain policeman who shows sadistical tendencies. I do not want to unravel the quantificational details here.

F. Object left of subject, both within VP.

...[_{V P} NP_{j,obj} [_{V P} NP_{subj} [_{V P} PA [_{V P} t_{j}... V ]]]

8.22 ...weil einen/ Alpengipfel ein Linguist/ enthusiastisch/ besteigt\.

(because) one  mountainpeak a linguist  enthusiastically climbs

,,For one of the peaks in the Alpes, there is a linguist who enthusiastically climbs it“ . Astonishingly, in this case it is possible to understand the multi-accent as wide focus over the whole rest of the sentence, with thus getting an indefinite subject NP, and a partitive object NP.

8.23 ... weil einen/ Alpengipfel ein Linguist enthusiastisch\ besteigt.

,,One of the peaks in the Alpes that are climbed by a linguist, is climbed by the linguist enthusiastically“ . Here we find the possibility to incorporate the indefinite subject NP into the verbal predicate and thus having it in the restrictor of the quantification (i.e. in the background of the sentence). The example is somewhat marginal but changes in acceptability with the adverbial in question. So it is fine to say:

8.24 ... weil einen/ Alpengipfel ein Linguist mit dem Seil\ bestiegen hat.

(because) one  mountainpeaka linguist  with the rope climbed has

,,One of the peaks climbed by a linguist was climbed by the linguist with a rope“ . Here it is extremely bad, however, to put the adverbial phrase to the front as in

8.25 ... weil einen/ Alpengipfel mit dem Seil\ ein Linguist bestiegen hat.

and expect this sentence to have the same meaning as 8.24. With other adverbials it is the other way round. For temporal phrases like ,,am Nachmittag“ (in the afternoon), the fronted version with incorporated subject indefinite is much better than the one with subject left to adverbial:

8.26 ... weil einen/ Alpengipfel am Nachmittag\ ein Linguist

(because) one  mountainpeak in the afternoon  a linguist

bestiegen hat

climbed has

,,One of the peaks climbed by linguistst was climbed by the linguist in the afternoon“

8.27 ... weil einen/ Alpengipfel ein Linguist am Nachmittag\ bestiegen hat

- here the subject ,,ein Linguist“ can only be read partitive or generic, but never indefinite. The latter case: subject NP generic, is already an example for the last structure I want to exemplify.

G. Object outside, subject outside VP, object left of subject

...[_{S} NP_{i,obj} [_{S} NP_{j,subj} [_{V P} GEN [_{V P} t_{j} [_{V P} PA [_{V P} t_{i}... V ]]]

8.28 ...weil einen/ Alpengipfel ein Linguist enthusiastisch\ besteigt.

,,One of the mountain peaks in the Alpes that are climbed by linguists is climbed enthusiastically by the typical linguist.“ Partitive object NP, generic subject NP. When Nps once are outside VP the word order seems to be comparatively free; at least with respect to the possibility to spread partitive and generic readings. -

As I promised above, I am sure that I missed quite a lot of intonation patterns and readings, and probably even some which would be relevant in the context. I have not discussed scope relations in any of the examples given here. I left this out on purpose, because at least in examples with raise-fall intonation patterns scope relations are interdependent with the two ways to interpret raise and fall, as I have scetched in section 8. So no comprehensive study of scope relations will be possible without a safe understanding of raise-fall intonation patterns and how they work. This is also a general feeling I got in investigating matters around the adverb in German. All these subjects, word order, intonation pattern, interpretation of NPs and interpretation and acceptability (as we have seen in section7) of adverbs of manner influence one another and can only be understood when investigated as a whole entangled complex.

9 Bibliography

Buering, D. (1993) ,,Word order in the German Middle Field and the Position of NegP.“ Manuscript, University of Koeln.

Buering, D. (1994) ,,Topic.“ manuscript, University of Koeln.

Cinque, G. (1993) ,,A Null Theory of Phrase and Compound Stress.“ Linguistic Inquiry 24, 239 - 297

Diesing, M. (1990) ,,The Syntactic Roots of Semantic Partition.“ Ph.D. dissertation, UMass, Amherst.

Diesing, M. (1992) Indefinites. Cambridge, Massachusetts: MIT Press.

von Fintel, K. (1994) ,,Restrictions on Quantifier Domains.“ Ph.D. dissertation, MIT.

Frey, W. (1993) Syntaktisch Bedingungen fuer die semantische Interpretation. Akademie Verlag: Berlin.

Geilfuß, J. (1992) ,,Nominal Quantifiers and Association with Focus“ , ConSole I proceedings, Sole Publications: Amsterdam.

Hamblin, C. L. (1973) ,,Questions in Montague English.“ Foundations of Language (1973), 41 - 53.

Heim, I. (1982) ,,The Semantics of Definite and Indefinite Noun Phrases.“ Ph.D. dissertation, UMass, Amherst.

Heim, I. (1990) ,,E-Type Pronouns and Donkey Anaphora.“ Linguistics and Philosophy 13, 137 - 177.

Herburger, E. (1991) ,,Focus and the Logical Form of Quantification.“ In Proceedings of the 1991 SALT Conference. Dept. of Modern Languages and Literatures, Cornell University.

Hoehle,T. N. (1982) ,,Explikationen fuer ,,normale Betonung“ und ,,normale Wortstellung“ .“ In Abraham, W. (ed.) Satzglieder im Deutschen, 75 - 153. Tuebingen: Narr.

Jaeger, G. (1994) ,,Topik, Fokus und schwache Quantoren.“ handout at GGS Tuebingen 1994, Arbeitsgruppe Strukturelle Grammatik, Berlin.

Kamp, J. A. W. (1981) ,,A Theory of Truth and Semantic Representation.“ In J. Groenendijk, T. Janssen, and M. Stokhof (eds.) Formal Methods in the Study of Language, 277 - 321. Amsterdam: Mathematical Centre.

Krifka, M. (1984) ,,Fokus, Topik, syntaktische Struktur und semantische Interpretation.“ manuscript, University of Muenchen.

Krifka, M. (1990) ,,Four thousand ships passed through the lock: object induced measure functions on events.“ Linguistics and Philosophy 12, 607 - 653.

Krifka, M. (1991) ,,A compositional Semantics for Multiple Focus Constructions.“ In Proceedings of the 1991 SALT Conference, 127 - 158. Dept. of Modern Languages and Literatures, Cornell University.

Milsark, G. (1974) ,,Existential Sentences in English.“ Ph.D. dissertation, MIT.

Muskens, R. (1991) ,, Meaning and Partiality.“, Ph.D. dissertation, University of Amsterdam.

Partee, B. H. (1988) ,,Many Quantifiers.“ In Proceedings of ESCOL 1988. Department of Linguistics, Ohio State University.

Partee, B. H. (1991) ,,Topic, Focus, and Quantification.“ In Proceedings of the 1991 SALT Conference. Dept. of Modern Languages and Literatures, Cornell University.

Rooth, M. E. (1985) ,,Association with Focus.“ Ph.D. disstertation, UMass, Amherst.

Rooth, M. E. (1992) ,,A Theory of Focus Interpretation.“ Natural Language Semantics 1, 75 - 116

von Stechow, A. (1989) ,,Focusing and Backgrounding Operators.“ Technical Report 6, Fachgruppe Sprachwissenschaft, University of Konstanz.

Westerståhl, D. (1985) ,,Determiners and context sets.“ In van Benthem, J. and ter Meulen, A. (eds.) Generalized quantifiers in natural language, 47 - 71. Dordrecht: Foris.